### v2

# Topological remarks on end and edge-end spaces

Abstract
The notion of ends in an infinite graph $G$ might be modified if we consider them as equivalence classes of infinitely edge-connected rays, rather than equivalence classes of infinitely (vertex-)connected ones. This alternative definition yields to the edge-end space $\Omega_{E}(G)$ of $G$, in which we can endow a natural (edge-)end topology. For every graph $G$, this paper proves that $\Omega_{E}(G)$ is homeomorphic to $\Omega(H)$ for some possibly another graph $H$, where $\Omega(H)$ denotes its usual end space. However, we also show that the converse statement does not hold: there is a graph $H$ such that $\Omega(H)$ is not homeomorphic to $\Omega_{E}(G)$ for any other graph $G$. In other words, as a main result, we conclude that the class of topological spaces $\Omega_{E}=\{\Omega_{E}(G):G\text{ graph}\}$ is strictly contained in $\Omega=\{\Omega(H):H\text{ graph}\}$.
1 Introduction:
In a graph $G$, a ray is an one-way infinite path, whose infinite connected subgraphs are called its tails. Similarly, a two-way infinite path is referred as a double-ray. Over the set $\mathcal{R}(G)$ of rays of $G$, one can define the two (similar) equivalence relations below:
  • End relation ($\sim$): For $r,s\in\mathcal{R}(G)$, we write $r\sim s$ whenever $r$ and $s$ are infinitely connected, i.e., there are infinitely many (vertex-)disjoint paths connecting $r$ and $s$. Equivalently, no finite set $S\subset V(G)$ separates $r$ and $s$, in the sense that the tails of $r$ and $s$ belong to different connected components of $G\setminus S$. Then, $\sim$ is indeed an equivalence relation on $\mathcal{R}(G)$, whose equivalence class of a ray $r$ (written as $[r]$) is called an end of $G$. The quotient $\mathcal{R}(G)/\sim$, referred as the end space of $G$, is denoted by $\Omega(G)$;
  • Edge-end relation ($\sim_{E}$): Now, for $r,s\in\mathcal{R}(G)$, we write $r\sim_{E}s$ whenever $r$ and $s$ are infinitely edge-connected, i.e., there is an infinite family of edge-disjoint paths connecting infinitely many vertices of $r$ to infinitely many vertices of $s$. Equivalently, no finite set $F\subset E(G)$ separates $r$ and $s$, in the sense that the tails of $r$ and $s$ belong to different connected components of $G\setminus F$. Under $\sim_{E}$, the equivalence class of a ray $r$ is denoted by $[r]_{E}$ and is called its edge-end. Then, we fix the notation $\Omega_{E}(G)=\mathcal{R}(G)/\sim_{E}$ for the edge-end space of $G$.
Roughly speaking, $\sim_{E}$ is obtained from $\sim$ after changing the roles of vertices by edges. In particular, $\sim_{E}$ is a weaker identification, since infinitely (vertex-)connected rays are also infinitely edge-connected. If $G$ is locally finite, then both equivalence relations turn out to be the same. However, when there are vertices of infinite degree, $\sim_{E}$ might identify more rays. For example, the graph presented by Figure 1 has two ends but only one edge-end.
$v_{-3}$$v_{-2}$$v_{-1}$$v_{0}$$v_{1}$$v_{2}$$v_{3}$$v_{\infty}$
Figure 1: This graph has two ends, corresponding to the rays $v_{0}v_{1}v_{2}\dots$ and $v_{0}v_{-1}v_{-2}\dots$, for example. However, by being infinitely edge-connected, it has only one edge-end.
In the literature, the definition of $\sim$ as just presented was first introduced by Halin in [9], although the notion of ends can be identified in previous algebraic discussions carried out by Hopf in [10] and by Freudenthal in [7]. Since then, a sort of results concerning finite graphs was generalized for infinite ones with the support of $\Omega(G)$. Regarding locally finite graphs, for example, the reader might consult the survey of Diestel in [5].
In its turn, inspired by that previous work of Halin, the definition of $\sim_{E}$ is due to Hahn, Laviolette and Širáň in [8]. Besides this paper, few other references in the literature mention the edge-end structure of infinite graphs, although it is naturally suggested by the standard end notion when considering an edge-related viewpoint. This incomplete state of art may be justified by the coincidence of $\sim$ and $\sim_{E}$ when restricted to locally finite graphs, so that the latter relation can be omitted in several studies within infinite graph theory. Despite that, the preprint [1] exemplifies how edge-ends might be suitable for discussing edge-connectivity results in arbitrary infinite graphs, even allowing, for instance, extensions of Menger-type theorems.
In any case, when introducing the quotient spaces $\Omega(G)$ and $\Omega_{E}(G)$, the corresponding topologies below arise from the separating notions displayed by vertices and edges:
  • End space topology: For an end $[r]\in\Omega(G)$, we fix a finite set $S\subset V(G)$ to define the following open basic neighborhood:
    $\Omega(S,[r])=\{[s]\in\Omega(G):S\text{ does not separate }r\text{ and }s\}.$
  • Edge-end space topology: For an edge-end $[r]_{E}\in\Omega_{E}(G)$, we fix a finite set $F\subset E(G)$ to define the following open basic neighborhood:
    $\Omega_{E}(F,[r]_{E})=\{[s]_{E}\in\Omega_{E}(G):F\text{ does not separate }r\text{ and }s\}.$
By highlighting topological properties of the above definitions, this paper aims to investigate their differences even if the underlying graphs are changed. More precisely, let us denote by $\Omega$ and $\Omega_{E}$ the classes of topological spaces that arise, respectively, as end spaces and edge-end spaces of graphs. By considering disjoint union of $|X|$ rays, we can quickly check that $\Omega$ and $\Omega_{E}$ contain every discrete space $X$, for example. Actually, $\Omega$ and $\Omega_{E}$ also contains all the complete ultrametric spaces, which are the end spaces of trees (see Theorem 3.1 of [15]). In fact, we will describe in Section 2 how every edge-end space can be obtained as the end space of a (possibly another) graph, proving that $\Omega_{E}\subseteq\Omega$. On a more challenging direction, the studies carried out by Section 4 concludes that the Alexandroff duplicate of the Cantor set is not an edge-end space of any graph, although Example 2.6 of [15] shows that this topological space belongs to $\Omega$. To summarize, Sections 2-4 prove the following main result:
Theorem 1.1.
$\Omega_{E}$ is a proper subclass of $\Omega$, i.e., every edge-end space is the end space of some graph, but the converse does not hold.
The proof of Theorem 1.1 relies on the techniques developed by Kurkofka and Pitz in [13]. In addition, we also explore their results aiming to study a topological game that, together with a special subbase, identifies when a topological space $X$ is actually an end space. This is done in a dialogue with the recent answer gave by Pitz in [15] to the problem of describing topologically the family $\Omega$, proposed by Diestel in [3]. Here, a special subbase for $X$ means a clopen one that is $\sigma-$disjoint, noetherian and nested, as introduced by Pitz in [15] and revisited in Section 5. The basis generated by a special subbase is thus called a special basis.
Then, we define the end game (denoted by $\mathrm{End}_{\mathcal{B}}$) over a zero-dimensional topological space $X$ with a fixed basis $\mathcal{B}$ to be the game where Players $\mathrm{I}$ and $\mathrm{II}$ alternate their moves according to the following rules:
  • The Player $\mathrm{I}$ starts the game by declaring a basic open set $U_{0}\subset X$ and the Player $\mathrm{II}$ answers with an open cover $\mathcal{U}_{0}$ for $U_{0}$ composed by pairwise disjoint basic open sets;
  • At the $n-$th round of the game, with $n\geq 1$, Player $\mathrm{I}$ declares a basic open set $U_{n}$ that is contained in some $V_{n-1}\in\mathcal{U}_{n-1}$. After, Player $\mathrm{II}$ answers with an open cover $\mathcal{U}_{n}$ for $U_{n}$ whose elements are pairwise disjoint basic open sets.
The assumption that $X$ is a zero-dimensional space guarantees that Player $\mathrm{II}$ has always a possible answer for moves of Player $\mathrm{I}$. If, at the end of the match, there are uniques $x\in X$ and $A\subset X\setminus\{x\}$ an open set such that $\bigcap_{n\in\mathbb{N}}U_{n}=\{x\}\cup A$, we declare Player $\mathrm{II}$ as the winner of the game. Hence, Player $\mathrm{I}$ wins the match otherwise. If Player $\mathrm{II}$ has a (stationary) wining strategy for $\mathrm{End}_{\mathcal{B}}$, we denote $\mathrm{II}\uparrow\mathrm{End}_{\mathcal{B}}(X)$. Similarly, $\mathrm{I}\uparrow\mathrm{End}_{\mathcal{B}}(X)$ means that Player $\mathrm{I}$ has a (stationary) winning strategy. Within these definitions, in Section 5 we also prove the following result:
Theorem 1.2.
Let $X$ be a Hausdorff topological space. Then, $\mathrm{II}\uparrow\mathrm{End}_{\mathcal{B}}(X)$ for some special basis $\mathcal{B}$ of $X$ if, and only if, $X$ is the end space of some graph.
2 Edge-end spaces as certain end spaces
In this section we prove the first part of Theorem 1.1, by including the class of edge-end spaces into the class of end spaces. To that aim, it is useful to identify when a vertex is infinitely connected to a ray. Hence, we say that $v\in V(G)$ dominates a ray $r$ if, for every finite set $S\subset V(G)\setminus\{v\}$, the tail of $r$ lies on the same connected component of $v$ in $G\setminus S$ Clearly, this is equivalent to the existence of an infinite family of paths connecting $v$ and $r$ that are disjoint unless by $v$. Therefore, $v$ dominates $r$ if, and only if, it dominates any ray equivalent to $r$, allowing us to say that $v$ dominates the end $[r]$. On the other hand, if $v$ and a tail of $r$ belong to different connected components of $G\setminus S$, we say that the finite set $S\subset V(G)\setminus S$ separates $v$ and $r$, as well as $v$ from $[r]$.
Rewriting the above definitions in an edge-related viewpoint, we say that $v\in V(G)$ edge-dominates the ray $r$ if, for every finite set $F\subset E(G)$, the vertex $v$ and a tail of $r$ belong to the same connected component of $G\setminus F$. Equivalently, there is an infinite family of edge-disjoint paths connecting $v$ and to infinitely many vertices of $r$. Hence, $v$ edge-dominates $r$ if, and only if, it dominates any other representative of the edge-end $[r]_{E}$. In particular, $v$ also edge-dominates every representative from $[r]$, in which case we say that $v$ edge-dominates the end $[r]$.
On the other hand, if $v$ and a tail of $r$ belong to different connected components of $G\setminus F$, we say that some finite set $F\subset E(G)$ separates $v$ and $r$, or even $v$ from $[r]_{E}$. Within these definitions, the fact that $\Omega_{E}\subseteq\Omega$ follows by an argument of connectedness:
Theorem 2.1.
Let $G$ be a graph. Then, there is a graph $H$ whose vertices edge-dominates at most one end (as defined by the relation $\sim$) and such that $\Omega(H)\cong\Omega_{E}(G)$.
Throughout this section, we will prove the above result by considering $H$ a graph obtained from $G$ after expanding some vertices to cliques. Formally, for each vertex $v\in V(G)$ that edge-dominates a ray, we denote by $K_{v}$ a complete graph of order $d(v)$, the degree of $v$. Hence, it is possible to fix $\rho_{v}:N(v)\to V(K_{v})$ a bijection between $N(v):=\{u\in V(G):uv\in E(G)\}$ and the vertices of $K_{v}$. Writting by $D$ the set of vertices of $G$ that edge-dominate some ray, the vertex set of $H$ may be given by $V(H)=(V(G)\setminus D)\cup\displaystyle\bigcup_{v\in D}V(K_{v})$, so that the elements of the (disjoint) union $\bigcup_{v\in D}V(K_{v})$ are called expanded vertices. The edge set of $H$, in its turn, is given by the (disjoint) union $\bigcup_{v\in D}E(K_{v})$ and the range of the function $\rho:E(G)\hookrightarrow E(H)$ defined as follows:
  • $\rho(uv)=uv$ if $u,v\in V(G)\setminus D$;
  • $\rho(uv)=u\rho_{v}(u)$ if $u\in V(G)\setminus D$ but $v\in D$;
  • $\rho(uv)=\rho_{u}(v)\rho_{v}(u)$ if $u,v\in D$.
Roughly speaking, $\rho$ is a map that attaches different edges incident in a vertex $v\in D$ to different vertices of $K_{v}$. Since adjacencies between vertices of $V(G)\setminus D$ are preserved, we regard $\rho$ as an inclusion map from $E(G)$ to $E(H)$. For a vertex $x\in K_{v}$, hence, we define its canonical edge to be the unique edge of $\rho(E(G))$ that has $x$ as an endpoint.
This inclusion map also translate rays of $G$ into rays of $H$ in a natural sense. In fact, if $r=v_{0}v_{1}v_{2}v_{3}\dots$ is a ray in $G$, we consider the ray $\rho(r)$ in $H$ whose presentation by its edges is $\rho(v_{0}v_{1})s_{1}\rho(v_{1}v_{2})s_{2}\rho(v_{2}v_{3})\dots$, where
  • $s_{i}=\emptyset$ is the empty edge if $v_{i}\notin D$, for $i\geq 1$;
  • $s_{i}=\rho_{v_{i}}(v_{i-1})\rho_{v_{i}}(v_{i+1})$ is the edge in $K_{v_{i}}$ connecting the canonical edges $\rho(v_{i-1}v_{i})$ and $\rho(v_{i}v_{i+1})$, for $i\geq 1$.
Clearly, finite paths of $G$ can be recovered in $H$ by the same construction. In other words, if $P=v_{0}v_{1}\dots v_{n}$ is a path in $G$, we denote by $\rho(P)$ the path in $H$ whose edges are $\rho(v_{0}v_{1})s_{1}\rho(v_{1}v_{2})s_{2}\rho(v_{2}v_{3})\dots s_{n-1}\rho(v_{n-1}v_{n})$. Conversely, a path or a ray $r$ in $H$ has the form $\rho(s)$ for some path or some ray $s$ of $G$, accordingly, whenever $|E(r)\cap E(K_{v})|\leq 1$ for every $v\in D$. In this sense, via $\rho$, the next technical result allows us to map edge-disjoint families of paths in $G$ to disjoint families of paths in $H$:
Lemma 2.2.
Fix $r$ a ray in $G$. Let $R$ be a vertex or a ray that cannot be separated from $r$ by finitely many edges. Then, there is an infinite family $\{P_{n}\}_{n\in\mathbb{N}}$ of edge-disjoint paths such that:
  1. 1.
    For every $n\in\mathbb{N}$, $P_{n}$ connects $r$ and $R$, i.e., it has one endpoint in $r$ and the other in $R$;
  2. 2.
    If $n\neq m$, then $V(P_{n})\cap V(P_{m})\subset D$.
Proof.
We first remark that, if $R$ is a vertex, then our main hypothesis actually says that $R$ dominates $r$. In this case, if another vertex $v\in V(G)$ cannot be separated from $R$ by finitely many edges, then $v\in D$ as well. This because $v$, $R$ and a tail of $r$ belong to the same connected component of $G\setminus F$, for every finite set $F\subset E(G)$.
Then, let $P_{0}$ be a path connecting $r$ and $s$. For instance, for some $n\geq 1$ suppose that we have so far defined finitely many edge-disjoint paths $P_{0},P_{1},\dots,P_{n-1}$ that connect $r$ and $R$. Moreover, we assume by induction that $V(P_{i})\cap V(P_{j})\subset D$ whenever $i\neq j$. But, by the observation made in the previous paragraph, any vertex from $V(G)\setminus D$ can be separated from $r$ and $R$ by a finite set of edges. In particular, there is a finite set $F\subset E(G)$ that separates every vertex of $\bigcup_{i=0}^{n-1}V(P_{i})\setminus D$ from $r$ and $R$. By hypothesis, in $G\setminus F$ there is a path $P_{n}$ that connects the tail of $r$ to $R$ (if it is a vertex) or to its tail (if it is a ray). Moreover, $V(P_{n})\cap V(P_{i})\subset D$ for every $i<n$ by the choice of $F$.
At the end of this recursive process, $\{P_{n}\}_{n\in\mathbb{N}}$ is the claimed family of paths. ∎
In particular, if $R=s$ is a ray that is edge-equivalent to $r$, then $\{P_{n}\}_{n\in\mathbb{N}}$ as in the above statement is a family of edge-disjoint paths connecting $r$ and $s$ whose elements intersect possibly at dominating vertices. Hence, in $H$, the family $\{\rho(P_{n})\}_{n\in\mathbb{N}}$ turns out to be a family of vertex-disjoint paths connecting $\rho(r)$ ro $\rho(s)$, since every expanded vertex is an endpoint of precisely one edge. This means that the map
$\begin{array}[]{cccc}\Phi:&\Omega_{E}(G)&\to&\Omega(H)\\ &[r]_{E}&\mapsto&[\rho(r)]\end{array}$
is well-defined. Proving Theorem 2.1, our aim now is to verify that $\Phi$ is an homeomorphism. First, we remark that its surjection also follows from Lemma 2.2:
Proposition 2.3.
The map $\Phi$ is surjective.
Proof.
Fix $r=x_{0}x_{1}x_{2}\dots$ a ray in $H$. First, consider the case in which $V(r)\cap V(K_{v})$ is finite for every $v\in D$. We will recursively define a ray $r^{\prime}=x_{0}^{\prime}x_{1}^{\prime}x_{2}^{\prime}\dots$ of $H$ as follows:
  • We declare $x_{0}^{\prime}=x_{0}$. If $x_{0}\in V(K_{v})$ for some $v\in D$, we also define $x_{1}^{\prime}=x_{i_{1}}$, in which $i_{1}=\max\{j\in\mathbb{N}:x_{j}\in V(K_{v})\}$. Since $K_{v}$ is a complete graph, $x_{0}^{\prime}$ and $x_{1}^{\prime}$ are indeed neighbors;
  • We suppose that $x_{0}^{\prime},x_{1}^{\prime},x_{2}^{\prime},\dots,x_{n}^{\prime}$ are already defined for some $n\in\mathbb{N}$. By induction, we can write $x_{k}^{\prime}=x_{i_{k}}$ for every $1\leq k\leq n$, for certain indices $i_{1}<i_{2}<i_{3}<\dots<i_{n}$. Moreover, we assume that, if $x_{n}^{\prime}\in V(K_{v})$ for some $v\in D$, then $i_{n}=\max\{j\in\mathbb{N}:x_{j}\in V(K_{v})\}$. We hence set $x_{n+1}^{\prime}=x_{i_{n}+1}$. In addition, if $x_{i_{n}+1}\in V(K_{v})$ for some $v\in D$, we also set $x_{n+2}^{\prime}=x_{i_{n+2}}$, where $i_{n+2}=\max\{j\in\mathbb{N}:x_{j}\in V(K_{v})\}$.
Defined this way, the ray $r^{\prime}$ is actually a subgraph of $r$, so that $[r^{\prime}]=[r]$. By construction, $|E(r)\cap E(K_{v})|\leq 1$ for every $v\in D$, so that $r^{\prime}=\rho(s)$ for some ray $s$ in $H$. Therefore, $\Phi([s]_{E})=[r^{\prime}]=[r]$.
Now, suppose that $V(r)\cap V(K_{v})$ is infinite for some $v\in D$. By definition of $D$, there is $s$ a ray in $G$ that $v$ edge-dominates. Then, there is $\{P_{n}\}_{n\in\mathbb{N}}$ an infinite family of edge-disjoint paths connecting $v$ and $s$. By Lemma 2.2, we can assume that $P_{n}\cap P_{m}\subset D$ if $n\neq m$. Since expanded vertices are endpoints of precisely one canonical edge, $\{\rho(P_{n})\}_{n\in\mathbb{N}}$ is a family of vertex-disjoint paths that verifies the equivalence between $\rho(s)$ and a ray $r_{v}$ composed by expanded vertices of $K_{v}$. On the other hand, since $V(r)\cap V(K_{v})$ is infinite and $K_{v}$ is a complete graph, we have $r\sim r_{v}$. Hence, $\Phi([s]_{E})=[\rho(s)]=[r_{v}]=[r]$.
When showing how $\Phi$ is injective, we can actually conclude that distinct edge-ends of $G$ are mapped to ends of $H$ that are not infinitely edge-connected:
Proposition 2.4.
If $r$ and $s$ are rays of $G$ such that $[r]_{E}\neq[s]_{E}$, then $\rho(r)$ and $\rho(s)$ are not edge-equivalent in $H$. In particular, $[\rho(r)]\neq[\rho(s)]$.
Proof.
Fix $r$ and $s$ two rays that can be separated by a finite set $F\subset E(G)$. In particular, $\rho(F)$ is also finite. For instance, suppose that, in $H\setminus\rho(F)$, there is a path $P=x_{0}x_{1}\dots x_{n}$ connecting the tails of $\rho(s)$ and $\rho(r)$. Suppose that $|V(P)|$ is minimum with that property. Then, given distinct $x_{i},x_{j}\in K_{v}$ for some $v\in D$, we must have $j=i+1$ if $i<j$, because the subpath $P^{\prime}=x_{0}x_{1}\dots x_{i}x_{j}\dots x_{n}$ is well defined and also connects $\rho(s)$ to $\rho(r)$. Hence, if $x_{i}\in D$, the edges $x_{i-1}x_{i}$ and $x_{i+1}x_{i+2}$, if exist, are canonical. In other words, $P=\rho(Q)$ for some path $Q$ that connects $r$ and $s$ in $G$. However, this contradicts the fact that $F$ separates the rays $r$ and $s$, because $E(P)\subset E(H)\setminus\rho(F)$ and, thus, $E(Q)\subset E(G)\setminus F$. Therefore, $\rho(F)$ separates $\rho(r)$ and $\rho(s)$. ∎
Corollary 2.5.
Every vertex of $H$ edge-dominates at most one end of $\Omega(H)$.
Proof.
Fix $v\in V(H)$ a vertex that edge-dominates two distinct ends of $H$. Fix $r$ and $s$ representatives for those ends. In particular, if $F\subset E(H)$ is finite, there is $P$ a path in $H\setminus F$ connecting $v$ to a tail of $r$. Then, $F^{\prime}=F\cup E(P)$ is also a finite set of edges, so that there is a path $P^{\prime}$ in $H\setminus F^{\prime}$ connecting $v$ to a tail of $s$. Therefore, by concatenating $P$ and $P^{\prime}$, we obtain a path in $H\setminus F$ connecting the tails of $r$ and $s$. This proves that $r$ and $s$ are infinitely edge-connected, although belong to different ends of $H$. Since $\Phi$ is surjective, this contradicts Proposition 2.4. ∎
Proving Theorem 2.1, then, we finish this section by showing that $\Phi$ is open and continuous:
Proposition 2.6.
$\Phi$ is an homeomorphism.
Proof.
We will first verify that $\Phi$ is continuous. To this aim, let $\Omega(S,[r])$ be a basic open set in the end space of $H$, for some $[r]\in\Omega(H)$ and some finite $S\subset V(H)$. Since $\Phi$ is a bijection, we can write the representative $r$ as $r=\rho(s)$ for some ray $s$ in $G$. If a fixed vertex $u\in S$ belongs to $D$, let $F_{u}=\{f_{u}\}$ denote the singleton set containing its canonical edge $f_{u}$. If not, $u$ does not edge-dominate the ray $s$, so that there is a finite set $F_{u}\subset E(G)$ that separates $s$ and $u$. Hence, $F=\displaystyle\bigcup_{u\in S}F_{u}$ is a finite set of edges. We then claim that the basic open set $\Omega_{E}(F,[s]_{E})$ for the edge-end space of $G$ is contained in $\Phi^{-1}(A)$. In fact, if $F$ does not separate $s$ and a ray $s^{\prime}$, there is $P$ a path connecting $s$ and $s^{\prime}$ in $G\setminus F$. By the choice of $F$, therefore, $\rho(P)$ is a path connecting $\rho(s)$ and $\rho(s^{\prime})$ in $G\setminus S$. In other words, $S$ does not separate $\rho(s)$ and $\rho(s^{\prime})$, proving that $\Phi(\Omega_{E}(F,[s]_{E}))\subset\Omega(S,[r])$.
Conversely, in order to show that $\Phi$ is an open map, let $\Omega_{E}(F,[s]_{E})$ be a basic open set containing an edge-end $[s]_{E}\in\Omega_{E}(G)$, for some finite $F\subset E(G)$. Hence, it is also finite the set $S=\{x\in V(H):x\text{ is endpoint of }\rho(e)\text{ for some }e\in F\}$. Then, it is enough to verify that $\Omega(S,[\rho(s)])\subset\Phi(\Omega(F,[s]_{E}))$. To this aim, again by the fact that $\Phi$ is surjective, an element $[r]\in\Omega(S,[\rho(s)])$ has a representative of the form $r=\rho(s^{\prime})$ for some ray $s^{\prime}$ of $G$. Since $S$ does not separate $\rho(s^{\prime})$ and $\rho(s)$, there is a path $P$ connecting the tails of these two rays in $H\setminus S$. As in Proposition 2.4, if we consider $P$ to have as few vertices as possible, we can write $P=\rho(Q)$ for some path $Q$ in $G\setminus F$ connecting $s$ and $s^{\prime}$. Hence, $F$ does not separate $s$ and $s^{\prime}$, so that $[s^{\prime}]_{E}\in\Omega_{E}(F,[s]_{E})$ and, therefore, $[r]=[\rho(s^{\prime})]\in\Phi(\Omega_{E}(F,[s]_{E}))$. ∎
3 Construction of edge-end spaces
Throughout this section, we will prove the converse of Theorem 2.1. Although this is not needed to show that $\Omega_{E}$ is a proper subclass of $\Omega$, Theorem 2.1 and the below result combined give a graph-theoretic description of edge-end spaces:
Theorem 3.1.
Let $G$ be a graph in which every vertex edge-dominates at most one end of $\Omega(G)$. Then, there is a graph $H$ such that $\Omega_{E}(H)\simeq\Omega(G)$.
We start the construction of $H$ by enveloping pre-described sets of ends of $G$, a procedure recently developed by Kurkofka and Pitz in [13]. In our setting, however, we will need only the restricted version below:
Lemma 3.2 ([13], Theorem 3.2).
Let $\varepsilon\in\Omega(G)$ be an end of $G$ and fix $\mathcal{R}_{\varepsilon}\subset\varepsilon$ a maximal family of pairwise (vertex-)disjoint rays. Denote by $D_{\varepsilon}$ the set of vertices that dominate $\varepsilon$. Then,
$E_{\varepsilon}=D_{\varepsilon}\cup\displaystyle\bigcup_{r\in\mathcal{R}_{\varepsilon}}V(r)$ (1)
has finite adesion, i.e., for each connected component $C$ of $G\setminus E_{\varepsilon}$, the set $N(C)=\{v\in E_{\omega}:v\text{ has a neighbor in }C\}$ is finite. Moreover, if $r$ is a ray such that $V(r)\cap E_{\varepsilon}$ is infinite, then $[r]=\varepsilon$.
Proof.
Fix $C$ a connected component of $G\setminus E_{\varepsilon}$. Let $\hat{C}$ be a subgraph of $G$ containing $C$, $N(C)$ and precisely one edge connecting each vertex of $N(C)$ to (some vertex of) $C$. In particular, $\hat{C}$ is also a connected subgraph of $G$. For instance, suppose that $N(C)$ is infinite, so that, by the Star-Comb Lemma (see [6]), one of the following assertions holds:
  • There is a star in $\hat{C}$ centered at some vertex $v\in V(C)$ with teethes in $N(C)$. In other words, for every finite set $S\subset V(G)\setminus\{v\}$, there is a path in $\hat{C}\setminus S$ connecting $v$ to a vertex of $N(C)$;
  • There is a comb in $\hat{C}$ with teethes in $N(C)$. More precisely, there is a ray $s$ in $C$ such that, for every finite set $S\subset V(G)$, there is a path connecting a vertex of $N(C)$ to the tail of $s$ in $\hat{C}\setminus S$.
Suppose that the former case occur and let $v\in V(C)$ be the center of the mentioned star. Fix a finite set $S\subset V(G)\setminus\{v\}$. Since the family $\mathcal{R}_{\varepsilon}$ is composed by pairwise disjoint rays, $S$ meets only finitely many of them. Hence, by adding initial segments of these rays, there is a finite set $S^{\prime}\subset V(G)\setminus\{v\}$ containing $S$ such that $r\setminus S^{\prime}$ is the tail of $r$ for every $r\in\mathcal{R}_{\varepsilon}$. As in the first item above, however, there is in $V(G)\setminus S^{\prime}$ a path $P$ connecting $v$ to a vertex $u$ of $N(C)$. If $u\in V(r)$ for some $r\in\mathcal{R}_{\varepsilon}$, then $u$ belongs to a tail of $r$ by our choice of $S^{\prime}$. If $u\in D_{\varepsilon}$, there is a path $P$ in $G\setminus S^{\prime}$ connecting $u$ to a tail of a ray whose end is $\varepsilon$, since $u$ dominates $\varepsilon$. In any case, $S$ does not separate $v$ and $\varepsilon$, contradicting the fact that $D_{\varepsilon}\cap V(C)=\emptyset$.
Suppose now that there is a ray $s$ in $C$ as in the second item above. Again, let $S\subset V(G)$ be any finite set and, as argued in the previous paragraph, assume that $r\setminus S$ is the tail of $r$ in $G\setminus S$ for every $r\in\mathcal{R}_{\varepsilon}$. By our main hypothesis over $s$, there is a path in $G\setminus S$ that connects its tail to a vertex $u\in N(C)$. If $u\in r$ to some $r\in\mathcal{R}_{\varepsilon}$, then $u$ belongs to a tail of $s$. If not, $u\in D_{\varepsilon}$, so that there is a path in $G\setminus S$ connecting $u$ to a tail of some ray whose end is $\varepsilon$. In any case, $S$ does not separate $s$ from $\varepsilon$, so that $[s]=\varepsilon$. However, since $s$ is contained in $C$, this contradicts the maximality of the family $\mathcal{R}_{\varepsilon}$.
Finally, let $r$ be a ray such that $V(r)\cap E_{\varepsilon}$ is infinite. Then, if $S\subset V(G)$ is finite, there is $v\in V(r)\cap E_{\varepsilon}\setminus S$. If $v$ dominates $\varepsilon$, then there is a path in $G\setminus P$ connecting $v$ to $r^{\prime}$ for some $r^{\prime}\in\varepsilon$. If not, then $v$ itself belongs to a ray $r^{\prime}\in\varepsilon$. In any case, $S$ does not separate $r$ from representatives of $\varepsilon$, proving that $[r]=\varepsilon$. ∎
Then, for each end $\varepsilon\in\Omega(G)$, we will fix $E_{\varepsilon}$ as defined by (1) and will call it an envelope for $\varepsilon$. By our main hypothesis over $G$, each vertex $v\in V(G)$ edge-dominates at most one such end $\varepsilon\in\Omega(G)$. If this is the case, we define a new vertex $v^{\prime}$ and a bipartition $\tau_{v}:N(v)\to\{v,v^{\prime}\}$ according to the following rules:
[…]
From now on in this section, $D\subset V(G)$ will denote the set of vertices that edge-dominates an end of $G$. Then, we consider $V(G)\cup\{v^{\prime}:v\in D\}$ as the vertex set of the claimed graph $H$. Its edge set is $\{vv^{\prime}:v\in D\}\cup\tau(E(H))$, in which $\tau:E(G)\hookrightarrow E(H)$ is the injective map given by:
[…]
In other words, $H$ is built from $G$ after duplicating vertices of $D$, rearranging the edges of $G$ according to $\tau$ and defining an edge between a vertex and its copy. In particular, by tracking the edges of $G$, the map $\tau$ also allows us to include paths and rays of $G$ into $H$. For example, given a ray $r=v_{0}v_{1}v_{2}\dots$ in $G$, we denote by $\tau(r)$ the ray in $H$ whose presentation by its edges is $\tau(v_{0}v_{1})s_{1}\tau(v_{1}v_{2})s_{2}\tau(v_{2}v_{3})s_{3}\dots$, in which, for $i\geq 1$:
  • $s_{i}=\emptyset$ is the empty edge if $v_{i}\notin D$ or $v_{i}\in D$ and $\tau_{v_{i}}(v_{i+1})=\tau_{v_{i}}(v_{i-1})$;
  • $s_{i}=v_{i}v_{i}^{\prime}$ if $v_{i}\in D$ and $\tau_{v_{i}}(v_{i+1})\neq\tau_{v_{i}}(v_{i-1})$.
Regarding the above notation, a path $P=v_{0}v_{1}v_{2}\dots v_{n}$ in $G$ can also be seen within $H$, via the path $\tau(P)$ whose edges are given by $\tau(v_{0}v_{1})s_{1}\tau(v_{1}v_{2})s_{2}\tau(v_{2}v_{3})s_{3}\dots s_{n-1}\tau(v_{n-1}v_{n})$. In particular, if $\{P_{n}\}_{n\in\mathbb{N}}$ is an infinite family of (vertex-)disjoint paths connecting the rays $r$ and $s$ in $G$, then $\{\tau(P_{n})\}_{n\in\mathbb{N}}$ are (vertex-)disjoint paths connecting $\rho(r)$ and $\rho(s)$. Therefore, the map below is well-defined and it is a natural candidate to be an homeomorphism between $\Omega(G)$ and $\Omega_{E}(H)$:
$\begin{array}[]{cccc}\Psi:&\Omega(G)&\to&\Omega_{E}(H)\\ &[r]&\mapsto&[\tau(r)]_{E}\end{array}$
In order to verify that $\Psi$ is indeed a bijection, the following technical observation is helpful:
Lemma 3.3.
Fix $K^{\prime}$ a connected subgraph of $H$ and let $K$ be the subgraph of $G$ whose edges are given by $\tau^{-1}(E(K^{\prime}))$ (and whose vertices are its endpoints). Then, $K$ is connected.
Proof.
We will prove that there is no empty cut in $K$. To that aim, let $\{A_{1},A_{2}\}$ be a bipartition of $V(K)$ into non-empty parts. For a contradiction, assume that there is no edge of $K$ with an endpoint in $A_{1}$ and the other in $A_{2}$. However, there are edges $e=u_{1}v_{1}$ and $f=u_{2}v_{2}$, for some $u_{1},v_{1}\in A_{1}$ and $u_{2},v_{2}\in A_{2}$. Since $K^{\prime}$ is connected, there is a path in this graph whose presentation via its edges is $e_{0}e_{1}e_{2}\dots e_{n}$, where $e_{0}=\tau(e)$ and $e_{n}=\tau(f)$. Consider the index
$i=\max\{0\leq j\leq n:e_{j}\in\tau(E(G))\text{ and }\tau^{-1}(e_{j})\text{ is an edge of }K[A_{1}]\}.$
This index is well-defined by the choice of $e_{0}=e$, but $i<n$ by the choice of $f=e_{n}$. Then, $e_{i+1}$ has one of the following forms, both contradicting the fact that $A_{1}\cap A_{2}=\emptyset$:
  • If $e_{i+1}=vv^{\prime}$ for some $v\in D$, then, by construction, $e_{i+2}\in\tau(E(G))$. Moreover, both $\tau^{-1}(e_{i})$ and $\tau^{-1}(e_{i+2})$ have $v$ as an endpoint. However, $\tau^{-1}(e_{i+2})$ is an edge of $K[A_{2}]$ by the maximality of $i$. Therefore, we should have $v\in A_{1}\cap A_{2}$;
  • If $e_{i+1}\in\tau(E(G))$, then $\tau^{-1}(e_{i+1})$ is an edge of $K[A_{2}]$ by the maximality of $i$. On the other hand, since $e_{i}\in\tau(E(G))$ is adjacent to $e_{i+1}$ in $H$, the edges $\tau^{-1}(e_{i})$ and $\tau^{-1}(e_{i+1})$ have a common endpoint $v\in V(G)$. Then, we should have again $v\in A_{1}\cap A_{2}$.
The fact that $\Psi$ is surjective can be seen as an application of König’s Lemma, while $\Psi$ is injective due to the main hypothesis over $G$:
Proposition 3.4.
$\Psi$ is bijective.
Proof.
Let $r^{\prime}$ be a ray in $H$, whose presentation via edges might be written as $f_{0}f_{1}f_{2}\dots$. Since the edges $\{vv^{\prime}:v\in D\}\subset E(H)$ are pairwise non-adjacent, $\tau^{-1}(E(r^{\prime}))$ is an infinite set of edges in $G$. Consider $K$ the subgraph of $G$ that contains precisely these edges and its endpoints, being connected by the above Lemma. We observe that every vertex $v\in V(K)$ has degree at most $4$. In fact, if $v\notin D$, then $v\in V(H)$ has at most two neighbors in $r^{\prime}$. Similarly, $v$ and $v^{\prime}$ have at most two neighbors each in $r^{\prime}$ if $v\in D$, so that $v$ is the endpoint of at most four edges in $K$ in this case. By König’s Lemma, then, there is $r$ a ray in $K$. Since $E(K)=\tau^{-1}(E(r^{\prime}))$, the ray $\tau(r)$ meets $r^{\prime}$ in infinitely many edges, so that $[\tau(r)]_{E}=[r^{\prime}]_{E}$. This verifies the surjection of $\Psi$.
In order to conclude that $\Psi$ is injective, let $s$ and $r$ non-equivalent rays in $G$. In other words, there is a finite set $S\subset V(G)$ that separates $r$ and $s$. Since $[s]\neq[r]$, by our main hypothesis over $G$ no vertex of $S$ edge-dominates both $r$ and $s$. Hence, for each $v\in S$ there is $F_{v}\subset E(G)$ finite that separates $v$ from $r$ or $s$. We claim that the finite set $\bigcup_{v\in S}\tau(F_{v})$ separates $\tau(r)$ and $\tau(s)$ in $H$. For instance, suppose that the tails of $\tau(s)$ and $\tau(r)$ belong to the same connected component $K^{\prime}$ of $H\setminus\displaystyle\bigcup_{v\in S}\tau(F_{v})$. Then, by Lemma 3.3, there is a path $P$ in $G$ connecting $r$ and $s$ such that $\tau(E(P))\subset H\setminus\displaystyle\bigcup_{v\in S}\tau(F_{v})$. However, there is $v\in V(P)\cap S$, because $S$ separates $r$ and $s$. Since $E(P)\cap F_{v}=\emptyset$, this contradicts the fact that $F_{v}$ separates the vertex $v$ from $r$ or $s$. ∎
The continuity of $\Psi$ follows easily from the fact that any edge-separator has a natural vertex-separator associated, an argument also employed in Proposition 2.6 when we verified that $\Phi$ is an open map:
Proposition 3.5.
$\Psi$ is continuous.
Proof.
Let $F\subset E(H)$ be a finite set of edges and fix $\eta\in\Omega_{E}(H)$ an edge-end of $H$. Since $\Psi$ is surjective, there is a ray $r$ in $G$ such that $[\tau(r)]_{E}=\eta$. For each $e\in F\cap\tau(E(G))$, consider $S_{e}$ the set of endpoints of $\tau^{-1}(e)$. If $e\in F$ has the form $e=vv^{\prime}$ for some $v\in D$, define $S_{e}=\{v\}$. Then, $S=\displaystyle\bigcup_{e\in F}S_{e}$ is finite.
Let $C$ be the connected component of $G\setminus S$ in which there is a tail of $r$. If $s$ is any other ray in $C$ and $P$ is a path connecting $r$ and $s$, then $\tau(P)$ connects $\tau(r)$ to $\tau(s)$ in $H\setminus F$ by construction. This argument verifies the inclusion $\Psi(\Omega(S,[r]))\subset\Omega_{E}(F,\eta)$, proving that $\Psi$ is continuous. ∎
We observe that the criteria for defining the neighbors of $v$ and $v^{\prime}$ in $H$, whether $v\in D$, was not mentioned in the proofs of the previous propositions. In fact, finishing this section, it is employed only to show that $\Psi$ is an open map:
Proposition 3.6.
$\Psi$ is an open map.
Proof.
Let $S\subset V(G)$ be finite and fix $C$ the connected component of $G\setminus S$ in which there is a ray $s$. Denote $\varepsilon=[s]$ and recall that we fixed a set $E_{\varepsilon}\subset V(G)$ as in (1) in order to define $\tau$. Choosing the representative $s$ so that $V(s)\subset E_{\varepsilon}$, we will now show that $\Psi(\Omega(S,[s]))$ is open in $\Omega_{E}(H)$.
First, for each $v\in S\setminus E_{\varepsilon}$, we observe that there is $C_{v}$ a connected component in $G\setminus E_{\varepsilon}$ containing $v$. By Lemma 3.2, the set $N(C_{v})=\{u\in E_{\varepsilon}:u\text{ has a neighbor in }C_{v}\}$ is finite. Moreover, for each $u\in N(C_{v})$, one of the options below is verified:
  • If $u\in D$, define the singleton set $F_{u}=\{uu^{\prime}\}\subset E(H)$;
  • If $u\notin D$, there is a finite set $F_{u}^{\prime}\subset E(G)$ that separates $u$ from $r$. Hence, we define $F_{u}=\tau(F_{u}^{\prime})$.
In any case, $F_{1}=\displaystyle\bigcup_{v\in S\setminus E_{\varepsilon}}\bigcup_{u\in N(C_{v})}F_{u}$ is a finite set of edges of $H$. Relying on the following claim, a similar set can be defined:
Claim: It is finite the set
$S^{\prime}=\{v\in E_{\varepsilon}:\text{there is no path connecting }v\text{ and a tail of }s\text{ in }G\setminus S\}.$
Proof of the Claim.
By definition, we first observe that $S$ separates vertices of $S^{\prime}\setminus S$ from $s$. Therefore, $D_{\varepsilon}\cap S^{\prime}\subset S$, where $D_{\varepsilon}$, as in (1), is the set of vertices of $G$ that dominate $\varepsilon$. If we assume, for a contradiction, that $S^{\prime}$ is infinite, then $S^{\prime}\cap\displaystyle\bigcup_{r\in\mathcal{R}_{\varepsilon}}V(r)$ is infinite, where $\mathcal{R}_{\varepsilon}$ is also defined by Lemma 3.2. If $S^{\prime}\cap V(r)$ is infinite for some $r\in\mathcal{R}_{\varepsilon}$, then there is a path connecting $s$ to a vertex of $S^{\prime}\cap V(r)$ in $G\setminus S$, since $[s]=\varepsilon=[r]$. This contradicts the definition of $S^{\prime}$, so that we may have $S^{\prime}\cap V(r)$ finite for every $r\in\mathcal{R}_{\varepsilon}$. In particular, $\mathcal{R}_{\varepsilon}^{\prime}=\{r\in\mathcal{R}_{\varepsilon}:S^{\prime}\cap V(r)\neq\emptyset\}$ is infinite and composed by pairwise disjoint rays. Hence, there is also some ray $r\in\mathcal{R}_{\varepsilon}^{\prime}$ such that $S\cap V(r)=\emptyset$, because $S$ is finite. Therefore, due to the equivalence between the rays $r$ and $s$, there is in $G\setminus S$ a path connecting a tail of $s$ to a vertex of $S^{\prime}\cap V(r)$. This contradicts the definition of $S^{\prime}$ once more, so that $S^{\prime}$ must be infinite. ∎
If $v\in S^{\prime}$ edge-dominates $s$, we define the set $J_{v}=\{vv^{\prime}\}\subset E(H)$. Otherwise, there is a finite set $J_{v}^{\prime}\subset E(G)$ that separates $s$ and $v$. In this case, we denote $J_{v}=\tau(J_{v}^{\prime})$. Then, $F_{2}=\displaystyle\bigcup_{v\in S^{\prime}}J_{v}$ is a finite set of edges of $H$.
Let $C^{\prime}$ be the connected component of $H\setminus F$ in which $\tau(s)$ has a tail. If another ray in that component has the form $\tau(s^{\prime})$ for some ray $s^{\prime}$ in $G$, then, by Lemma 3.3, there is a path $P$ in $G$ connecting $s$ and $s^{\prime}$ such that $\tau(E(P))\subset E(H)\setminus F$.
On the other hand, if $[s]\neq[s^{\prime}]$, by Lemma 3.2 there is a connected component $C_{s^{\prime}}$ in $G\setminus E_{\varepsilon}$ in which $s^{\prime}$ has a tail. Hence, writing $P$ in terms of its vertices as $v_{0}v_{1}v_{2}\dots v_{n}$, also assuming that $v_{0}\in E_{\varepsilon}$ and $v_{n}\in C_{s^{\prime}}$, fix $i=\min\{0\leq j\leq n:v_{j}\in E_{\omega}\text{ and }v_{j+1}\in C_{s^{\prime}}\}$. Since $v_{i}\in E_{\varepsilon}$, one of the following cases must hold:
  • If $v_{i}$ edge-dominates $s$, the edge $v_{i}v_{i}^{\prime}$ is defined in $H$. Moreover, by definition of $\tau_{v_{i}}$, the path $P$ must contain this edge, because $v_{i+1}\in C_{s^{\prime}}$. Hence, $v_{i}v_{i}^{\prime}$ does not belong to $F$. By definition of $F_{1}$, this means that $S\cap C_{s^{\prime}}=\emptyset$, while, by definition of $F_{2}$, $v_{i}\notin S^{\prime}$;
  • Supposing now that $v_{i}$ can be separated from $S$ by finitely many edges, we have $S\cap C_{s^{\prime}}=\emptyset$. Otherwise, $E(P)\cap F_{v_{i}}^{\prime}\neq\emptyset$, because $P$ connects the tails of $s$ and $s^{\prime}$ and contains $v_{i}$, contradicting the fact that $\tau(E(P))\subset H\setminus F_{1}$. Analogously, if $v\in S^{\prime}$ for instance, then we have $E(P)\cap J_{v_{i}^{\prime}}\neq\emptyset$ by the same reason, contradicting the fact that $\tau(E(P))\subset H\setminus F_{2}$.
In both cases, we conclude that $S\cap C_{s^{\prime}}=\emptyset$ and $v\notin S^{\prime}$. Then, $C_{s^{\prime}}$ is a connected subgraph of $G\setminus S$ containing $v_{i+1}$ and a tail of $s^{\prime}$, while there is also a path in $G\setminus S$ connecting $v_{i}$ and a tail of $s$. Therefore, the tails of $s$ and $s^{\prime}$ belong to the same connected component of $G\setminus S$, proving that $\Omega_{E}(F,[s]_{E})\subset\Psi(\Omega(S,[s]))$. ∎
4 Topological implications of Theorem 2.1
Although Theorems 2.1 and 3.1 present a characterization of edge-end spaces in terms of the end spaces as originally defined in the literature, it is not immediate to assess whether these classes of topological spaces are the same. Actually, edge-end spaces define a strict smaller subfamily of all the end spaces, which this Section aims to conclude in order to finish the proof of Theorem 1.1.
However, this investigation relies on the work of Kurkofka and Pitz in [13], where these authors proved that end-spaces of graphs are precisely the topological spaces arising from ray spaces of special (order) trees. In particular, given a graph $G$, they constructed a tree $T$ such that the end-space of $G$ is also the end-space of some graphs on $T$. As the program carried out by this Section, we will verify that, if $G$ is under the hypothesis of Theorem 3.1, such construction can lead to a countable order tree $T$, assuming convenient topological properties. Since countable graphs have metrizable end-spaces, by admitting normal (graph-theoretic) spanning trees for instance, this will prove the following:
Theorem 4.1.
Let $X$ be a Lindelöf and first-countable topological space. If $X$ is the edge-end space of some graph, then $X$ is metrizable.
Introducing the main notation of this section, we recall that a tree, in a set-theoretic context, is a partially ordered set $\langle T,\leq\rangle$ where $\mathring{\lceil t\rceil}=\{s\in T:s<t\}$ is well-ordered by $\leq$ for every node $t\in T$. The order type $\mathrm{h}(t)$ of $\mathring{\lceil t\rceil}$ is the height of the node $t\in T$, so that, for a given ordinal $\alpha$, the set $\mathcal{L}_{\alpha}(T)=\{t\in T:t\text{ has height }\alpha\}$ is the $\alpha$ level of $T$. If there is a predecessor $s=\max\mathring{\lceil t\rceil}$, we say that $t\in T$ is a successor point. Otherwise, the cofinality of $\mathring{\lceil t\rceil}$ is infinite and we say that $t$ is a limit point of $T$. For the least ordinal $\alpha$ such that $\mathcal{L}_{\alpha}(T)=\emptyset$, we say that $\alpha$ is the height of $T$.
Throughout this paper, trees will always be rooted, i.e., such that $|\mathcal{L}_{0}(T)|=1$. Moreover, a set $R\subset T$ in which the order $\leq$ is total is called a chain of $T$, while a maximal one is called a branch. If $T$ has no infinite branches, we say that $T$ is rayless. Following the notation of [13], we say that $R\subset T$ is a high-ray if it is a down-closed chain of cofinality $\omega$. Then, the tops of this high-ray are the minimal elements greater then every node of $R$.
Graph-theoretic trees, i.e., the acyclic connected graphs, are examples of trees whose height is bounded by $\omega$, if endowed with the usual tree-order after fixing a root. On the other hand, Brochet and Diestel in [2] introduced canonical graphs associated to a fixed order tree $T$. Following their terms, a $T-$graph is a graph whose vertex set is $T$ and such that, for every $t\in T$, its neighborhood $N(t)$ is cofinal in $\mathring{\lceil t\rceil}$. Moreover, if, for every limit node $t\in T$, there is a finite subset $X\subset\mathring{\lceil r\rceil}$ such that every $s>t$ has its neighbors below $t$ as elements of $X$, we say that $G$ is a uniform $T-$graph.
In particular, the endpoints of any edge in a $T-$graph are comparable in the order of $T$. Therefore, this definition generalizes the concept of normal spanning trees, whose first studies by Jung for infinite graphs (see [11] for example) provided useful tools for studying topological and combinatorial behaviour of ends. In this context, we recall that, for a given graph $G$ and a fixed subgraph $T$, a $T-$path in $G$ is a path that has precisely its endpoints in $H$. If $T$ is an order tree, we say that $T$ is normal if every $T-$path has comparable endpoints in the tree order.
Both in finite and infinite settings, normal trees in graphs are usually found by depth-search algorithms. For quite broad graph classes in the literature, such as for countable ones, these algorithms lead even to normal spanning trees. As an example to be recovered further in this Section, we outline the unified method developed by Pitz in [14] to obtain these structures:
Proposition 4.2 ([14], Theorem 3).
Let $G$ be any graph and fix a finite subset $K\subset V(G)$. Write $G^{\prime}$ for the graph $G\setminus K$. Then, $G^{\prime}$ has a maximal normal (graph-theoretic) tree $T$ such that, for every connected component $C$ of $G^{\prime}\setminus T$:
  • The neighborhood $N(C)=\{v\in G\setminus C:v\text{ has a neighbor in }C\}$ is infinite. In particular, $N(C)\cap T$ is contained in a infinite branch $r_{C}$ of $T$;
  • Every $v\in N(C)$ dominates $r_{C}$.
Proof.
We will construct an increasing sequence $T_{0}\subseteq T_{1}\subseteq T_{2}\subseteq T_{3}\subseteq\dots$ of rayless normal trees in $G^{\prime}$, by first fixing $T_{0}=\{r\}$ an arbitrary root. Suppose that $T_{n}$ is defined for some $n<\omega$. Let $\mathcal{C}_{n}$ denote the collection of the connected components of $G^{\prime}\setminus T_{n}$. Since $T_{n}$ is rayless and normal, the neighborhood $N(C_{n})=\{v\in G\setminus C:v\text{ has a neighbor }x\in C_{n}\}$ is finite for every $C_{n}\in\mathcal{C}_{n}$, because $K$ is finite and $N(C)\cap T_{n}$ is contained in a branch of $T$.
For every vertex $v\in N(C_{n})$, fix $x_{v}^{n}\in C_{n}$ one of its neighbors. Then, by the previous observation, $F_{C_{n}}=\{x_{v}^{n}:v\in N(C_{n})\}$ is finite. Moreover, if $r$ is a ray in $G$ that has a tail in $C_{n}\in\mathcal{C}_{n}$, then $N(C_{n})\cup F_{C_{n}}$ is a finite set of vertices that separates $r$ from any vertex of $D_{n}:=\displaystyle\bigcup_{C_{n}\in\mathcal{C}_{n}}F_{C_{n}}$. Due to this property, we say that $D_{n}$ is disperse. Hence, by a well-known result of Jung in [11] (see also Theorem 2.2 of [12]), there is [^1] For a more immediate construction of $T_{n+1}$, we may apply a routine depth-search algorithms within each connected component $C_{n}\in\mathcal{C}_{n}$, reaching the finitely many vertices of $F_{C_{n}}$. See Proposition 1.5.6 of [6] for details. $T_{n+1}$ a rayless normal tree in $G^{\prime}$ extending $T_{n}$ and containing $D_{n}$. Moreover, $T_{n+1}\cap C_{n}$ is connected for every $C_{n}\in\mathcal{C}$. After all, a path in $C_{n}$ connecting two vertices $v,u\in T_{n+1}\cap C_{n}$ must intersect $T_{n+1}$ in a third vertex if $u$ and $v$ are not comparable in its tree order, because $C_{n}\subset G^{\prime}\setminus T_{n}$ and $T_{n+1}$ is normal.
Once finished this inductive process, we claim that $T=\displaystyle\bigcup_{n<\omega}T_{n}$ satisfies the statement. We first observe that $T$ is a normal tree, because so is $T_{n}$ for each $n<\omega$. For instance, suppose that $N(C)$ is finite for some $C\in\mathcal{C}$, where $\mathcal{C}$ denotes the family of connected components of $G^{\prime}\setminus T$. Therefore, we must have $N(C)\cap T\subset T_{n}$ for some $n<\omega$. In this case, $C$ is also a connected component of $G^{\prime}\setminus T_{n}$, contradicting the fact that $F_{C}\subset T_{n+1}\setminus T_{n}$. Hence, the first item of the Proposition holds.
Now, consider $T^{\prime}\supsetneq T$ a (graph-theoretic) tree in $G^{\prime}$ extending the tree order of $T$, if it exists. Let $v\in T^{\prime}\setminus T$ be minimal, i.e., such that $\mathring{\lceil v\rceil}\subset T$. Note that $\mathring{\lceil v\rceil}$ is finite, due to the fact that $T$ has height bounded by $\omega$. Then, in $G^{\prime}\setminus\mathring{\lceil v\rceil}$ there is a path connecting $v$ to a vertex of $T$, because the connected component $C\in\mathcal{C}$ containing $v$ has infinitely many neighbors in $T$. Since $T^{\prime}$ extends the tree order of $T$, this verifies that $T^{\prime}$ is not normal. In other words, $T$ is a maximal normal (graph-theoretic) tree.
Finally, fixed $C\in\mathcal{C}$ and $v\in N(C)$, denote by $r_{C}$ the branch of $T$ containing $N(C)$. For a finite set $S\subset V(G)\setminus\{v\}$, let $n<\omega$ be big enough so that $v\in T_{n}$ and $S\cap T=S\cap T_{n}$. Consider $C_{n}\in\mathcal{C}_{n}$ the connected component of $G^{\prime}\setminus T_{n}$ in which $C\subset C_{n}$ and, hence, $v\in N(C_{n})$. In $T_{n+1}$, then, there is a path connecting $x_{v}^{n}$, a neighbor of $v$ in $C_{n}$, to a vertex $v^{\prime}$ of $r_{C}$, because $T_{n+1}\cap C_{n}\subset G^{\prime}\setminus S$ is connected. Therefore, $v$ dominates the ray $r_{C}$. ∎
Relying on the above Proposition, the proof of Theorem 4.1 is done by revisiting Theorem 1.1 of Kurkofka and Pitz in [13]. For a given graph $G$, their result is a decomposition of $V(G)$ in the shape of an order tree $T$, codifying $\Omega(G)$ through the high-rays of $T$. More precisely, a pair $(T,\mathcal{V})$ is called a partition tree of $G$ if $\mathcal{V}=\{V_{t}:t\in T\}$ is a partition of $V(G)$ into connected subsets satisfying the properties below:
  • $|V_{t}|=1$ if $t\in T$ is not a limit point;
  • The graph $\dot{G}:=\frac{G}{\mathcal{V}}$ obtaining by contracting the parts of $\mathcal{V}$ to single vertices is a $T-$graph;
  • For each successor $t\in T$, the neighborhood
    $N(V_{\lfloor t\rfloor})=\left\{u\in V(G)\setminus\displaystyle V_{\lfloor t\rfloor}:u\text{ has a neighbor in }\displaystyle V_{\lfloor t\rfloor}\right\}$
    is finite, where $V_{\lfloor t\rfloor}=\displaystyle\bigcup_{s\geq t}V_{s}$. In this case, we say that $(T,\mathcal{V})$ has finite adhesion.
Hence, by tracking some rays of $G$, we are able to describe high-rays of $T$. In other words, given $[r]\in\Omega(G)$ an end, it is well-defined the set
$\Theta([r])=\{t\in T:r\text{ has a tail in }\displaystyle V_{\lfloor t\rfloor}\}.$ (2)
This is clearly a down-closed chain of $T$ and, by Lemma 6.2 of [13], it has countable cofinality. If this cofinality is infinite, we say that $\Theta([r])$ corresponds to the end $[r]$, because $\Theta([r])$ is then an element of $\mathcal{R}(T)=\{\text{high-rays of }T\}$. When every $[r]\in\Omega(G)$ corresponds to precisely one high-ray of $T$, in the sense that $\Theta:\Omega(G)\to\mathcal{R}(T)$ is a bijection, we even say that the partition tree $(T,\mathcal{V})$ displays all the ends of $G$. The existence of partition trees with that property is guaranteed by Theorem 1.1 of [13], whose proof is partially adapted below to provide a convenient statement for the graph classes we aim to approach:
Theorem 4.3 ([13], Theorem 1.1).
Let $G$ be a connected graph whose vertices dominate at most one end. Suppose that $\Omega(G)$ is a first-countable Lindelöf topological space. Then, $G$ has a partition tree $(T,\mathcal{V})$ that display all its ends and such that:
  1. 1.
    $T$ has countable height, bounded by $\omega\cdot\omega$;
  2. 2.
    The subtree $\hat{T}=\{t\in T:t\text{ belongs to a high-ray of }T\}$ is countable.
Proof of Theorem 1.1 of [13] revisited.
For some cardinal $\kappa\leq\omega$, we will recursively construct a sequence of partition trees $\{(T_{n},\mathcal{V}_{n})\}_{n<\kappa}$ for $G$. Let $T_{0}^{\prime}$ be a maximal normal tree for the graph as in Proposition 4.2, whose tree-order is denoted by $\leq$. Consider $F_{0}$ the set of connected components of $G\setminus T_{0}^{\prime}$, so that each $C\in F_{0}$ has its infinite neighborhood $N(C)$ contained in a branch $r_{C}$ of $T_{0}^{\prime}$. Then, the order $\leq$ can be extended to $T_{0}:=T_{0}^{\prime}\cup F_{0}$, by declaring $C>t$ for every $C\in F_{0}$ and every $t\in r_{C}$. If we set $V_{t}^{0}=\{t\}$ for every $t\in T_{0}^{\prime}$ and $V_{C}^{0}=V(C)$ for every $C\in F_{0}$, we partition $G$ with the family $\mathcal{V}_{0}=\{V_{t}^{0}:t\in T_{0}^{\prime}\}\cup\{V_{C}^{0}:C\in F_{0}\}$. It is easily verified that $(T_{0},\mathcal{V}_{0})$ is a partition tree for $G$.
Now, for some $n<\omega$, suppose that we have defined a partition tree $(T_{n},\mathcal{V}_{n})$ of height bounded by $\omega\cdot(n+1)+1$. If the height of $T_{n}$ is precisely $\omega\cdot(n+1)$, we set $\kappa=n+1$ and finish the construction. Otherwise, by induction, the nodes of height $\omega\cdot n$ in $T_{n}$ define a set $F_{n}$ with the following form:
$C\in F_{n}\iff C\text{ is a connected component of }G\setminus\bigcup_{t\in T_{n}\setminus F_{n}}V_{t}.$ (3)
Moreover, for each $C\in F_{n}$, we assume that its neighborhood $N(C)=\{v\in V(G)\setminus C:v\text{ has a neighbor in }C\}$ is contained in a ray $r_{C}$ such that $\Theta([r_{C}])$ is a high-ray of $T_{n}$. Hence, for every $C\in F_{n}$ we can apply Lemma 7.2 of [13] to obtain $U_{C}\subset V(C)$ a connected vertex set that encodes suitable topological properties. Among these, we mention:
Fact: If $\mathcal{D}(C)$ is the set of connected components of $C\setminus U_{C}$, the subgraph $D\in\mathcal{D}(C)$ has finite neighborhood in $G$. In other words, the set $N(D)=\{v\in G\setminus D:v\text{ has a neighbor in }D\}$ is finite.
Then, rooted at a vertex that has some neighbor in $U_{C}$, we can fix $T_{D}$ a normal tree for $D$, obtained when applying Proposition 4.2 to $G[D\cup N(D)]$ with $K=N(D)$. Hence, every connected component $D^{\prime}$ of $D\setminus T_{D}$ has infinitely many neighbors within a branch $r_{D^{\prime}}$ of $T_{D}$. Moreover, since $N(D)$ is finite, $N(D^{\prime})=\{v\in V(G)\setminus D^{\prime}:v\text{ has a neighbor in }D^{\prime}\}$ is concentrated in $r_{D^{\prime}}$. Then, writing $\mathcal{V}_{n}=\{V_{t}^{n}\}_{t\in T_{n}}$ and $\mathcal{V}_{n+1}=\{V_{t}^{n+1}\}_{t\in T_{n+1}}$, the partition tree $(T_{n+1},\mathcal{V}_{n+1})$ can be described as follows:
  • The tree $T_{n+1}$ extends the order tree $T_{n}$, by additionally containing the nodes from $T_{D}$ for every $C\in F_{n}$ and every $D\in\mathcal{D}(C)$. In this case, we set $t>C$ for each $t\in T_{D}$. Moreover, for every connected component $D^{\prime}$ of $D\setminus T_{D}$, we also see $D^{\prime}$ as a node of $T_{n+1}$, defining $D^{\prime}>t$ for every $t\in r_{D^{\prime}}$;
  • We set $V_{t}^{n+1}=V_{t}^{n}$ for every $t\in T_{n}\setminus F_{n}$. Given $C\in F_{n}$ and $D\in\mathcal{D}(C)$, however, we define $V_{C}^{n+1}=U_{C}$ and $V_{t}^{n+1}=\{t\}$ for every $t\in T_{D}$. Finally, we set $V_{D^{\prime}}^{n+1}=V(D^{\prime})$ for every connected component $D^{\prime}$ of $D\setminus T_{D}$.
Then, $|V_{t}^{n+1}|=1$ for every successor $t\in T_{n+1}$ and $\frac{G}{\mathcal{V}_{n+1}}$ is indeed a $T_{n+1}-$graph. Relying on the above Fact, the choice of $V_{t}^{n+1}$ for a limit node $t\in F_{n}$ guarantees that $(T_{n+1},\mathcal{V}_{n+1})$ has finite adhesion. If $(T_{n},\mathcal{V}_{n})$ is defined for every $n<\omega$, we set $\kappa=\omega$. Then, writing $\mathcal{V}=\{V_{t}\}_{t\in T}$, the requested partition tree $(T,\mathcal{V})$ arises from the following limit definition:
  • We set $T=\displaystyle\bigcup_{n<\kappa}T_{n}$, extending the order of $T_{n}$ for every $n<\kappa$;
  • If $\kappa=n+1<\omega$, we consider $\mathcal{V}=\mathcal{V}_{n}$, so that $(T,\mathcal{V})$ is the partition tree $(T_{n},\mathcal{V}_{n})$. Otherwise, given $t\in T$, we set $V_{t}=V_{t}^{n}$ for any $n>\min\{i<\omega:t\in T_{i}\}$. In this case, we claim that $\mathcal{V}=\{V_{t}\}_{t\in T}$ is a partition of $V(G)$. For instance, suppose that there is a vertex $v\in V(G)\setminus\displaystyle\bigcup_{t\in T}V_{t}$. Being $(T_{n},\mathcal{V}_{n})$ a partition tree for each $n<\omega$, we must have $v\in C_{n}$ for some $C_{n}\in F_{n}$. By considering a big enough $n_{0}<\omega$, let $u\in\displaystyle\bigcup_{t\in T_{n_{0}}}V_{t}^{n_{0}}$ be a neighbor of $v$. By the second item of Proposition 4.2, $u$ dominates the ray $r_{C_{n}}$ for every $n\geq n_{0}$. However, if $n>m\geq n_{0}$, the choice of $U_{C_{m}}$ guarantees that $r_{C_{n}}$ and $r_{C_{m}}$ can be separated by finitely many vertices. In other words, $u$ dominates infinitely many non-equivalent rays, which is a contradiction. Hence, $\mathcal{V}$ is indeed a partition of $G$.
The verification that $T$ displays the ends of $G$ is precisely the one given by [13]. Since $T$ has countable height, bounded by $\omega\cdot\omega$, it remains to show that the subtree $\hat{T}=\{t\in T:t\text{ belongs to a high-ray of }T\}$ is itself countable.
For instance, suppose that $\hat{T}$ is uncountable and fix
$\alpha=\min\{\xi<\omega\cdot\omega:\text{the }\alpha\text{ level of }\hat{T}\text{ is uncountable}\}.$
That ordinal exists because $\hat{T}$ has countable height, and we have $\alpha>0$ since $\hat{T}$ is rooted. If $\alpha$ is a successor ordinal, written as $\alpha=\beta+1$, let $t\in\mathcal{L}_{\beta}(\hat{T})$ be a predecessor of an uncountable family of nodes $\{t_{i}\}_{i<\omega_{1}}\subset\mathcal{L}_{\alpha}(\hat{T})$. Also relying on the previous Fact and by passing $\{t_{i}\}_{i<\omega}$ to another uncountable subsequence if necessary, we can assume that $N(V_{\lfloor t_{i}\rfloor})=N(V_{\lfloor t_{j}\rfloor})=:S$ for every $i,j<\omega_{1}$, since $h(t)$ is countable and $(T,\mathcal{V})$ has finite adhesion. Then, $\{\Omega(S,\varepsilon):\varepsilon\in\Omega(G)\}$ is an open cover for $\Omega(G)$ whose distinct elements are disjoint. However, by definition of $\hat{T}$, there is $r_{i}$ a ray in $G$ that has a tail in $V_{\lfloor t_{i}\rfloor}$, for each $i<\omega_{1}$. Hence, $S$ separates $r_{i}$ and $r_{j}$ if $i\neq j$. This means that $\{\Omega(S,[r_{i}]):i<\omega_{1}\}\subset\{\Omega(S,\varepsilon):\varepsilon\in\Omega(G)\}$ is an uncountable subfamily whose elements are pairwise disjoint, contradicting the assumption that $\Omega(G)$ is a Lindelöf topological space.
Therefore, $\alpha$ must be a limit ordinal, so that $\mathcal{L}_{\alpha}(T)=F_{n}$ for some $n<\omega$. We argue that $n\neq 0$. Otherwise, fix $\{C_{i}\}_{i<\omega_{1}}\subset F_{0}\cap\hat{T}$ uncountable. Recall that these are connected components of $G\setminus T_{0}^{\prime}$, and, therefore, are pairwise disjoint. Then, one of the following cases is verified, but both lead to contradictions:
  • If there is an uncountable subset $I\subset\omega_{1}$ such that $r_{C_{i}}=r_{C_{j}}=:r$ for every $i,j\in I$, then each $C_{i}$ has infinitely many neighbors in the branch $r$ of $T_{0}^{\prime}$. As before, for each $i\in I$, let $r_{i}$ be a ray in $V_{\lfloor C_{i}\rfloor}$, whose existence is guaranteed by the definition of $\hat{T}$ and by the fact that $(T,\mathcal{V})$ displays the ends of $G$. Then, given a finite subset $S\subset V(G)$, we have $[r_{i}]\in\Omega(S,[r])$ for all but finitely many indices $i\in I$. This, however, contradicts the fact that $\Omega(G)$ is a first-countable topological space;
  • Then, there is an uncountable subset $I\subset\omega_{1}$ such that $r_{C_{i}}\neq r_{C_{j}}$ for every $i,j\in I$. In this case, $[r_{C_{i}}]\neq[r_{C_{j}}]$, because $(T,\mathcal{V})$ displays the ends of $G$. Since $r_{C_{i}}$ is a branch of $T_{0}^{\prime}$, we have $r_{C_{i}}\subset\hat{T}$ for each $i\in I$. Fix $v_{i}\in r_{C_{i}}$ a neighbor of the connected component $C_{i}$. Being $\hat{T}\cap T_{0}^{\prime}$ countable by the minimality of $\alpha$, there must be $v\in T_{0}^{\prime}$ such that $v=v_{i}$ for uncountably many indices $i\in I$. According to Proposition 4.2, this means that $v$ dominates uncountably many non-equivalent rays, contradicting the main hypothesis over $G$.
Then, we must have $n>0$. Moreover, $F_{n-1}\cap\hat{T}$ is countable, since this is a smaller (limit) level of $\hat{T}$. Hence, for some $C\in F_{n-1}\cap\hat{T}$, the set $\{t\in F_{n}\cap\hat{T}:t>C\}$ is uncountable. As an element of $\hat{T}$, the node $C$ has only countably many successors, because $\alpha$ is a limit ordinal. Therefore, we can fix $v_{0}\in\hat{T}$ a successor of $C$ such that $\{t\in F_{n}\cap\hat{T}:t>v_{0}\}$ is uncountable. By construction, we recall that $v_{0}$ is the root of a normal tree $T_{D}$ for a connected component $D$ of $C\setminus V_{C}^{t}$. Fixing an uncountable family $\{D_{i}^{\prime}\}_{i<\omega_{1}}\subset\{t\in F_{n}\cap\hat{T}:t>v_{0}\}$, then, each $D_{i}^{\prime}$ is a connected component of $D\setminus T_{D}$. Analogously to the above discussion, one of the following cases is verified, but both also lead to contradictions:
  • Suppose that there is an uncountable set $I\subset\omega_{1}$ such that $r_{D_{i}^{\prime}}=r_{D_{j}^{\prime}}=:r$ for every $i,j\in I$. Since $\{D_{i}^{\prime}\}_{i<\omega_{1}}\in F_{n}\cap\hat{T}$, the branch $r$ of $T_{D}$ is contained in $\hat{T}$. For each $i\in I$, let $r_{i}$ be a ray in $V_{\lfloor C_{i}\rfloor}$, whose existence is guaranteed by the definition of $\hat{T}$ and by the fact that $(T,\mathcal{V})$ displays the ends of $G$. Then, given a finite subset $S\subset V(G)$, we have $[r_{i}]\in\Omega(S,[r])$ for all but finitely many indices $i\in I$. This, however, contradicts the fact that $\Omega(G)$ is a first-countable topological space;
  • Then, there is an uncountable set $I\subset\omega_{1}$ such that $r_{D_{i}^{\prime}}\neq r_{D_{j}^{\prime}}$ for every $i,j\in I$. In this case, $[r_{D_{i}^{\prime}}]\neq[r_{D_{j}^{\prime}}]$, because $(T,\mathcal{V})$ displays the ends of $G$. Since $r_{D_{i}^{\prime}}$ is a branch of $T_{D}$ for every $i\in I$, we have $r_{D_{i}^{\prime}}\subset\hat{T}$, allowing us to choose $v_{i}\in\hat{T}\cap T_{D}$ a neighbor of $D_{i}^{\prime}$. Observing that $\hat{T}\cap T_{D}$ is countable by the minimality of $\alpha$, there is $v\in\hat{T}\cap T_{D}$ a vertex such that $v=v_{i}$ for every $i$ within some uncountable subset of $I$. By Proposition 4.2, this means that $v$ dominates uncountably many distinct ends of $G$, contradicting our main hypothesis over this graph.
Therefore, $\hat{T}$ is countable.
Actually, by the proof of Theorem 1.1 in [13], one can translate the convergence of ends in $G$ through combinatorial properties of the partition tree $(T,\mathcal{V})$ constructed above. As discussed in that paper, this is a sequentially faithful partition tree displaying all the ends of $G$, which encodes the topology of $\Omega(G)$ through the tree-structure of $T$. In our settings, Theorem 3 of [13] proves the following:
Proposition 4.4 ([13], Theorem 3).
Let $G$ be a graph as in Theorem 4.3. Then, there is $T^{\prime}$ an order tree such that $\Omega(G)$ is the end space of any uniform $T^{\prime}-$graph. Moreover, $T^{\prime}$ can be chosen so that $\hat{T^{\prime}}=\{t\in T^{\prime}:t\text{ belongs to a high-ray of }T^{\prime}\}$ is countable.
Proof of Theorem 3 of [13] revisited.
Fix the sequentially faithful partition tree $(T,\mathcal{V})$ for $G$ given by Theorem 4.3. Let $T^{\prime}$ be obtained from $T$ according to the procedure below:
  1. 1.
    For every limit node $t\in T$, denote by $S(t)$ the set of its successors, if there are some. Then, since $(T,\mathcal{V})$ has finite adhesion, the set $N_{s}:=N(V_{\lfloor s\rfloor})$ is finite for each $s\in S(t)$;
  2. 2.
    Now, for every limit node $t\in T$ that has a successor and every finite $X\subset\lceil t\rceil$, we declare a new node $v(t,X)$ to be a successor of $t$ and a predecessor of each $s\in S(t)$ with $N_{s}=X$. We then remove $t$.
Since only non-empty levels of $T$ were modified to construct $T^{\prime}$, the height of this latter tree is also countable. In addition, the proof that $\Omega(G)$ is the end space of any $T^{\prime}-$graph is precisely the one given in [13].
Finally, let $t^{\prime}\in T^{\prime}$ be a node that belongs to a high-ray of $T^{\prime}$. If $t^{\prime}$ is a successor node in $T^{\prime}$, then it is also a successor node in $T$ by construction of $T^{\prime}$. In particular, $t^{\prime}$ also belongs to a high-ray of $T$, so that $t^{\prime}\in\hat{T}$. If $t^{\prime}$ is a limit point, however, then $t^{\prime}=v(t,X)$ for some $t\in T$ that has at least one successor and some finite subset $X\subset\lceil t\rceil$. Actually, the node $t$ must lie on a high-ray of $T$ (i.e., $t\in\hat{T}$), because $t^{\prime}$ itself belongs to a high-ray of $T^{\prime}$. Hence, since $\hat{T}$ is countable and there are countably many finite subsets of $\lceil t\rceil$ for every $t\in\hat{T}$, it follows that $\hat{T^{\prime}}$ is also countable. ∎
In its original statement, Theorem 3 of [13] proves that every end space is the end space of some special order tree. For completeness, we recall that an order tree $T$ is special if it is a countable disjoint union of antichains, i.e., sets of pairwise incomparable elements. Then, there are two different ways to see $\mathcal{R}(T)$ as a topological space, called the ray space of the tree $T$. Formalized by Proposition 5.4 of [13], the first way consider $\mathcal{R}(T)$ as the end space of any uniform $T-$graph $G$, identifying $\mathcal{R}(T)$ and $\Omega(G)$ through the map $\Theta$ as in (2) after considering the partition tree $(T,\{t\}_{t\in T})$. The second way, introduced by Lemma 2.1 of [15], is defined intrinsically in terms of $T$, by declaring as basic open neighborhood around the high-ray $r\in\mathcal{R}(T)$ a set of the form
$[t,F]=\{s\in\mathcal{R}(T):t\in s\text{ and }t^{\prime}\notin s\text{ for every }t^{\prime}\in F\},$ (4)
where $t\in r$ and $F\subset T$ is a finite collection of tops of $r$. The equivalence between these two approaches is discussed by Max Pitz in subsection 2.4 of [15]. However, relying on both, we can conclude the main result of this section:
Theorem 4.5.
Let $G$ be a graph in which every vertex dominates at most one end. If $\Omega(G)$ is a Lindelöf first-countable topological space, then $\Omega(G)$ is metrizable.
Proof.
According to Proposition 4.4, there is an order tree $T$ such that $\Omega(G)$ is the end space of any uniform $T-$graph. Moreover, $T$ can be chosen so that the subtree $\hat{T}=\{t\in T:t\text{ belongs to a high-ray of }T\}$ is countable. However, $\mathcal{R}(T)$ and $\mathcal{R}(\hat{T})$ describe the same topological space, since the basic open neighborhoods given by (4) are coincident. In particular, if $G^{\prime}$ is a uniform $\hat{T}-$graph, its end space is homeomorphic to $\Omega(G)\simeq\mathcal{R}(T)$. However, $G^{\prime}$ is countable, because so is $\hat{T}$. Therefore, $\Omega(G^{\prime})$ is metrizable, since $G^{\prime}$ has a normal spanning tree (see [4] for instance). ∎
Combining Theorem 2.1 with this statement, Theorem 4.1 follows. As a consequence, we are ready to exhibit a topological space in $\Omega\setminus\Omega_{E}$, also finishing the proof of Theorem 1.1:
Corollary 4.6.
The Alexandroff duplicate of the Cantor set is not an edge-end space, although it is the end space of a graph.
Proof.
Let us revisit the Example 2.6 presented by Pitz in [15], that describes how the Alexandroff duplicate of the Cantor set can be seen as a ray space of a special tree $T$. To this aim, consider $T$ as obtained from the binary tree $2^{<\omega}$ after adding, for each branch $r\subset 2^{<\omega}$, a top $t_{r}$ and a countable increasing sequence of nodes $\{t_{r}=t_{r}^{0}<t_{r}^{1}<t_{r}^{2}<t_{r}^{3}<\dots\}$. We observe that $T$ is special, since it has height $\omega\cdot 2$ and its levels are antichains. Moreover, the space $\mathcal{R}(T)$ is first-countable, because the family of basic open neighborhoods described by 4.4 is countable. In particular, for each branch $r$ of $2^{<\omega}$, the high-ray $\overline{r}=r\cup\{t_{r}^{i}:i<\omega\}$ is an isolated point of $\mathcal{R}$. By Proposition 2.16 of [15], this is also a Lindelöf (actually compact) space, once the nodes of $T$ have countably many successors. Nevertheless, $\mathcal{R}(T)$ is not second-countable, because $\{\{\overline{r}\}:r\text{ is a branch of }2^{<\omega}\}$ is an uncountable family of pairwise disjoint open sets. Hence, $\mathcal{R}(T)$ is not a metric space. Therefore, in the graph-theoretical setting, Theorem 4.1 implies that the end space of any uniform $T-$graph [^2] In this case, a $T-$graph can be obtained by declaring successor nodes to be adjacent to their predecessors, while defining the edges between each top $t_{r}$ and the nodes of the corresponding branch $r\subset 2^{<\omega}$. For constructions of graphs on arbitrary special trees, see Theorem 4.6 of [13]. is not an edge-end space. ∎
5 End spaces via a topological game
As pointed out in the introduction, Diestel’s question regarding the topological characterization of end spaces was only recently solved by Pitz in [15]. In this section, we give an interpretation of his result via the topological game previously mentioned. This approach is reasonable since games codify some order trees, the objects that, under the appropriate conditions obtained by [13], represents end spaces faithfully.
We start this discussion by detailing the definition of the end game and explaining its motivation. As a big picture, Theorem 3 in [13] constructs an order tree that displays the topological behaviour of the ends in a given graph $G$. Our Theorem 1.2 approaches the dual direction: for a topological space $X$ with suitable hypothesis over a basis $\mathcal{B}$, we aim to define a special order tree $T$ whose high-ray space is $X$. The nodes of $T$ will be some open basic sets from $\mathcal{B}$, organized throughout $T$ so that a high-ray correspond to a $\subseteq$-⊆ -decreasing sequence $\{U_{n}\}_{n\in\mathbb{N}}\subset\mathcal{B}$. In this direction, regarding the basis $\mathcal{B}$, we recall that the end game $\mathrm{End}_{\mathcal{B}}$ between Players $\mathrm{I}$ and $\mathrm{II}$ is described by the following rules:
  • The Player $\mathrm{I}$ starts the game by declaring a basic open set $U_{0}\in\mathcal{B}$. Assuming that $X$ is zero-dimensional, as it is every end space, Player $\mathrm{II}$ is able to answer with an open cover $\mathcal{U}_{0}\subset\mathcal{B}$ for $U_{0}$ composed by pairwise disjoint elements;
  • At the $n-$th round of the game, with $n\geq 1$, Player $\mathrm{I}$ declares an open set $U_{n}\in\mathcal{B}$ that is contained in some $V_{n-1}\in\mathcal{U}_{n-1}$. As in the previous item, $\mathrm{II}$ answers with a pairwise disjoint family $\mathcal{U}_{n}$ of basic open sets that cover $U_{n}$.
Hence, at the end of the match, the moves of Player $\mathrm{I}$ describe a $\subseteq$-⊆ -sequence $\{U_{n}\}_{n\in\mathbb{N}}$. Therefore, we call $\mathrm{II}$ the winner of the game if this choice of moves captures a point $x\in X$ in the following sense: the intersection $\bigcap_{n\in\mathbb{N}}U_{n}$ can be uniquely written as a disjoint union $\{x\}\cup A$, where $A\subset X$ is an open set not containing $x$. Comparing this criteria with the open basic neighborhood in 4, $x$ will be further identified with a high-ray $r$ of some high-ray space, so that $A$ will correspond to the open set $\bigcup_{\begin{subarray}{c}t\text{ top }\\ \text{of }r\end{subarray}}[t,\emptyset]$.
In our context, a strategy for Player $\mathrm{I}$ is a map $\varphi$ that declares an answer $U_{n+1}=\varphi(\mathcal{U}_{0},\mathcal{U}_{1},\dots,\mathcal{U}_{n})$ for every sequence of moves $\{\mathcal{U}_{i}\}_{i=0}^{n}$ made by $\mathrm{II}$. Analogously, $\psi$ is a strategy for Player $\mathrm{II}$ if, for every sequence of moves $(U_{0},U_{1},\dots,U_{n})$ made by $\mathrm{I}$, the open cover $\mathcal{U}_{n}=\psi(U_{0},U_{1},\dots,U_{n})$ is an answer given by $\mathrm{II}$ at the $n-$th round of the game. Moreover, if $\psi$ depends only on the previous move of Player $\mathrm{I}$, we denote $\psi(U_{n})=\psi(U_{0},U_{1},\dots,U_{n})$ and call $\psi$ a stationary strategy. In this case, we say that $\psi$ is winning if $\mathrm{II}$ wins every match of the form
$\langle U_{0},\psi(U_{0}),U_{1},\psi(U_{1}),U_{2},\psi(U_{2}),\dots\rangle,$
where $\{U_{n}\}_{n\in\mathbb{N}}$ is the sequence of moves played by $\mathrm{I}$. When such a strategy exists, we denote $\mathrm{II}\uparrow\mathrm{End}_{\mathcal{B}}$. Similar definitions for stationary and winning strategies for Player $I$ can be settled, although we will not need them to our study. Actually, unless opposite mentions, all strategies in this paper are assumed to be stationary.
Intuitively, the winning criteria for $\mathrm{II}$ suggests that a winning strategy for this player is a map that displays a wide range of points in $X$. If $X$ is taken to be the end space of some graph, this motivates the characterization given by Theorem 1.2. Before proving it, we will precise some notations from its statement. Then, for a topological space $X$, we call a clopen subbase $\mathcal{C}=\{U_{\alpha}\}_{\alpha\in\Lambda}$ special if the following properties hold:
  • $\mathcal{C}$ is $\sigma-$disjoint, in the sense that it can be written as a countable union of antichains, i.e., families of pairwise disjoint open sets;
  • $\mathcal{C}$ is nested, namely, given $\alpha,\beta\in\Lambda$, then either $U_{\alpha}\cap U_{\beta}=\emptyset$, or $U_{\alpha}\subseteq U_{\beta}$ or $U_{\beta}\subseteq U_{\alpha}$;
  • $\mathcal{C}$ is noetherian, i.e., every $\subseteq$-⊆ -increasing chain of elements of $\mathcal{C}$ has a $\subseteq$-⊆ -maximum element.
If such a clopen subbase exists, we call $X$ a special topological space. Then, the end space characterization described by Pitz in [15] claims that $X$ is the end space of some graph if, and only if, it admits a clopen special subbase that is hereditarily complete. In our approach via topological games, this last hypothesis, to be further mentioned, will be replaced by the assumption that Player $\mathrm{II}$ wins when playing the end game in a special basis. Here, a special basis for $X$ means the basis generated by some clopen special subbase. In that direction, the following description is useful:
Lemma 5.1.
Let $X$ be a topological space and $\mathcal{C}=\{U_{\alpha}\}_{\alpha\in\Lambda}$ be a $\sigma-$disjoint, nested and noetherian collection of clopen sets. Then, $\mathcal{C}$ is a special subbase of $X$ if, and only if,
=$\mathcal{B}=\left\{U_{\alpha}\setminus\bigcup_{\beta\in F}U_{\beta}:U_{\alpha},U_{\beta}\in\mathcal{C},F\subset\Lambda\text{ finite}\right\}$ (5)
is a basis for $X$. In particular, the elements of $\mathcal{B}$ are also clopen sets.
Proof.
We will denote by $U^{c}:=X\setminus U$ the complement of a subset $U$ in $X$, also inspiring the notation $\mathcal{C}^{c}=\{U_{\alpha}^{c}\}_{{\alpha}\in\Lambda}$. Suppose first that $\mathcal{C}$ is a clopen subbase for $X$, i.e., that $\mathcal{C}\cup\mathcal{C}^{c}$ is an open subbase. Hence, $\left\{\displaystyle\bigcap_{\alpha\in F_{1}}U_{\alpha}\cap\bigcap_{\beta\in F_{2}}U_{\beta}^{c}:F_{1},F_{2}\subset\Lambda\text{ finite}\right\}$ is a basis for $X$. However, if $F_{1},F_{2}\subset\Lambda$ are finite and $\bigcap_{\alpha\in F_{1}}U_{\alpha}\neq\emptyset$, then $\{U_{\alpha}\}_{\alpha\in F_{1}}$ is totally ordered by $\subseteq$, since $\mathcal{C}$ is nested. Therefore, we can write $\bigcap_{\alpha\in F_{1}}U_{\alpha}\cap\bigcap_{\beta\in F_{2}}U_{\beta}^{c}=U_{\gamma}\cap\bigcap_{\beta\in F_{2}}U_{\beta}^{c}=U_{\gamma}\setminus\bigcup_{\beta\in F_{2}}U_{\beta},$ where $U_{\gamma}$ is the $\subseteq$-⊆ -minimum element from $\{U_{\alpha}\}_{\alpha\in F_{1}}$. Then, $\mathcal{B}$ is also a basis for $X$.
Conversely, suppose that $\mathcal{B}$ is a basis for $X$. For a basic open set $B=\displaystyle U_{\alpha}\setminus\bigcup_{\beta\in F}U_{\beta}\in\mathcal{B}$, we can write $U_{\alpha}\setminus\bigcup_{\beta\in F}U_{\beta}=U_{\alpha}\cap\bigcap_{\beta\in F}U_{\beta}^{c}$, so that $B$ is a finite intersection of elements from $\mathcal{C}\cup\mathcal{C}^{c}$. Therefore, $\mathcal{C}$ is a clopen subbase for $X$.
The basis description provided by 5 is quite similar to the basic open sets for ray spaces of trees in 4. Due to this, an open set of the form $U_{\alpha}\setminus\displaystyle\bigcup_{\beta\in F}U_{\beta}$ will now be denoted by $[U_{\alpha},F]$, for every $\alpha\in\Lambda$ and every finite $F\subset\Lambda$. Relying on this comparison, one can verify that $\mathcal{R}(T)$ is a special topological space for every special tree $T$:
Lemma 5.2.
If $T$ is a special tree, then $\mathcal{C}=\{[t,\emptyset]\}_{t\in T}$ is a special clopen subbase for $\mathcal{R}(T)$. In particular, the end space of any graph is a special topological space.
Proof.
We will first observe that, for any $t\in T$, the open set $[t,\emptyset]$ is also closed in $\mathcal{R}(T)$. Indeed, if $r\in\mathcal{R}(T)$ is a high-ray that does not contain $t$, then either $t\geq t_{0}$ for some top $t_{0}$ of $r$ or $s\in r$ for some node $s\in T$ incomparable with $t$. In the former case, $r\in[x,\{t_{0}\}]\subset\mathcal{R}(T)\setminus[t,\emptyset]$ for any $x\in r$, while, in the latter, $r\in[s,\emptyset]\subset\mathcal{R}(T)\setminus[t,\emptyset]$. In both case, it is verified that $[t,\emptyset]$ is a closed set of $\mathcal{R}(T)$. In other words, $\mathcal{C}$ is a family of clopen sets.
For nodes $s,t\in T$, we have the following relations:
  • If $s\leq t$, then every high-ray containing $t$ also contains $s$, so that $[t,\emptyset]\subset[s,\emptyset]$;
  • If $s$ and $t$ are incomparable, then no high-ray containing $s$ contains $t$, so that $[t,\emptyset]\cap[s,\emptyset]=\emptyset$.
In particular, $\mathcal{C}$ is a nested family. Moreover, if $\mathcal{D}=\{[t,\emptyset]\}_{t\in\Lambda}$ is a subset of $\mathcal{C}$ totally ordered by $\subseteq$, the above items show that $\Lambda\subset T$ is totally ordered as well. Then, there exists $t_{0}=\min\Lambda$, so that $[t,\emptyset]\subset[t_{0},\emptyset]$ for every $t\in T$. In other words, $\mathcal{D}$ is noetherian.
Since $T$ is special, we can write $T=\displaystyle\bigcup_{n\in\mathbb{N}}A_{n}$, where $A_{n}\subset T$ is an antichain for each $n\in\mathbb{N}$. By the second item above, $\mathcal{A}_{n}=\{[t,\emptyset]:t\in T\}$ is also an antichain in $\mathcal{R}(T)$. Therefore, $\mathcal{C}$ is $\sigma-$disjoint, since $\mathcal{C}=\displaystyle\bigcup_{n\in\mathbb{N}}\mathcal{A}_{n}$.
Finally, given $r\in\mathcal{R}(T)$, a node $t\in r$ and a finite set $F$ of tops of $r$, we have $[t,F]=[t,\emptyset]\setminus\displaystyle\bigcup_{s\in F}[s,\emptyset]$. In particular, the set
=$\mathcal{B}=\left\{[t,\emptyset]\setminus\displaystyle\bigcup_{s\in F}[s,\emptyset]:t\in T,F\subset T\text{ finite}\right\}$ (6)
is a basis for $\mathcal{R}(T)$. Therefore, $\mathcal{C}$ is a clopen subbase for $\mathcal{R}(T)$ by Lemma 5.1. ∎
Playing on the special basis described by 6, we can explicit a winning strategy for Player $\mathrm{II}$:
Proposition 5.3.
Let $T$ be a special tree. Then, Player $\mathrm{II}$ has a winning strategy for the end game in $\mathcal{R}(T)$ played in the special basis generated by $\{[t,\emptyset]\}_{t\in T}$.
Proof.
For every $t\in T$ we will denote by $\hat{t}$ the node $\hat{t}=\max\{s\in T:s\text{ is a limit node and }s\leq t\}$. In addition, if $t$ is a limit node, we will write $\{t_{i}\}_{i<\omega}$ for a fixed cofinal sequence in $\mathring{\lceil t\rceil}$. Let $\mathcal{B}$ denote the special basis generated by $\{[t,\emptyset]\}_{t\in T}$, as in the expression 6.
Within this notation, we are ready to explicit a winning (stationary) strategy for Player $\mathrm{II}$ in $\mathrm{End}_{\mathcal{B}}$. To this aim, fix a basic open set $V=[t,F]\in\mathcal{B}$. By definition of $[t,F]$ in 4, we can assume that $t<m$ for every $m\in F$. Then, consider the following three families of open sets:
  • For every $m\in F$, define the index $i(m)=\min\{i\in\mathbb{N}:t<\hat{m}_{i}\}$. For every successor $s$ of $t$, the basic open set $[s,\Lambda_{s}^{1}(F)]$ will be referred as a type $1$ set, where $\Lambda_{s}^{1}(F)=\{\hat{m}_{i(m)}:m\in F,\hat{m}_{i(m)}>s\}\cup F$;
  • For every $m\in F$, the open set $[\hat{m}_{i(m)},\Lambda_{m}^{2}(F)]$ will be referred as a type $2$ set, where
    $\Lambda_{m}^{2}(F)=(\{\hat{n}_{i(n)}:n\in F\}\cup\{\hat{n}:n\in F\}\cup F)\cap\{t\in T:t>\hat{m}_{i(m)}\};$
  • Finally, a type $3$ set is an open set of the form $[\hat{m},\Lambda_{m}^{3}(F)]$, where
    $\Lambda_{m}^{3}(F)=(\{\hat{n}_{i(n)}:n\in F\}\cup F)\cap\{t\in T:t>\hat{m}\}$
    for each $m\in F$.
Its is easily verified that the family
$\{[s,\Lambda_{s}^{1}(F)]:s\text{ is a successor of }t\}\cup\{[\hat{m}_{i(m)},\Lambda_{m}^{2}(F)]:m\in F\}\cup\{[\hat{m},\Lambda_{m}^{3}(F)]:m\in F\}$
covers $V=[t,F]$ with pairwise disjoint open sets. Hence, we will consider this open cover as the answer $\psi(V)$ of Player $\mathrm{II}$ when Player $\mathrm{I}$ declares $V$ in $\mathrm{End}_{\mathcal{B}}$. In order to see that $\psi$ is a winning strategy, let
$\langle V_{0},\psi(V_{0}),V_{1},\psi(V_{1}),V_{2},\psi(V_{2}),\dots\rangle$
be a match in which Player $\mathrm{II}$ follows the instructions of $\psi$. In particular, for each $n\in\mathbb{N}$, Player $\mathrm{I}$ chooses $V_{n+1}$ as an open set contained in $[t_{n},F_{n}]\in\psi(V_{n})$, for some $t_{n}\in T$ and some finite $F\subset T$. As in Lemma 5.2, $\{t_{n}\}_{n\in\mathbb{N}}$ is a totally ordered subset of $T$, because $[t_{n+1},F_{n+1}]\subset[t_{n},F_{n}]$ for every $n\in\mathbb{N}$. Hence, there is $r$ a high-ray of $T$ in which $\{t_{n}\}_{n<\omega}$ is a cofinal sequence.
In order to verify that Player $\mathrm{II}$ wins the match, we will prove that $\bigcap_{n\in\mathbb{N}}V_{n}$ can be written uniquely as a union of a high-ray and an open set not containing it. To that aim, we first observe that $\bigcap_{n\in\mathbb{N}}V_{n}=\displaystyle\bigcap_{n\in\mathbb{N}}[t_{n},F_{n}]$, because $V_{n+1}\subset[t_{n},F_{n}]\subset V_{n}$ for every $n\in\mathbb{N}$. Moreover, there is no open set containing $r$ that is contained in $\bigcap_{n\in\mathbb{N}}[t_{n},F_{n}]$, due to the fact that $\{t_{n}\}_{n\in\mathbb{N}}$ is cofinal in $r$. Hence, it is enough to prove that $\bigcap_{n\in\mathbb{N}}[t_{n},F_{n}]\setminus\{r\}$ is an open set in $\mathcal{R}(T)$.
Indeed, fix a high-ray $r^{\prime}\in\displaystyle\bigcap_{n\in\mathbb{N}}[t_{n},F_{n}]\setminus\{r\}$. Note that $r^{\prime}$ strictly contains $r$, once more because $\{t_{n}\}_{n\in\mathbb{N}}$ is cofinal in $r$. If $s$ denotes the top of $r$ belonging to $r^{\prime}$, we finish the proof by concluding that there is no $m>s$ with $m\in\displaystyle\bigcup_{n\in\mathbb{N}}F_{n}$. In particular, we must have $[s,\emptyset]\subset\displaystyle\bigcap_{n\in\mathbb{N}}[t_{n},F_{n}]$. For a contradiction, suppose that there is such a node $m>s$. If we choose $m$ with minimum height satisfying that property, we must have $m\in F_{n}$ for every $n$ bigger than some fixed $n_{0}\in\mathbb{N}$. In particular, for every $n\geq n_{0}$, the open set $[t_{n},F_{n}]$ is not of type $1$. Neither it is of type $3$: otherwise, $\hat{t_{n}}=\hat{m}$, so that the chain $\{t\in T:t_{n}\leq t\leq m\}$ is finite, contradicting the fact that $m>s$. Hence, the open set $[t_{n},F_{n}]$ is of type $2$ for every $n\geq n_{0}$. In particular, by the minimality of $m$, we must have $m=\hat{m}$ and the sequence $\{t_{n}\}_{n\geq n_{0}}$ must be cofinal in $\mathring{\lceil m\rceil}$. Therefore, $m=s$, contradicting the choice of $m$. ∎
From now on in this section, our aim is to prove a converse statement for Proposition 5.3. In other words, if $X$ is a Hausdorff topological space in which $\mathrm{II}\uparrow\mathrm{End}_{\mathcal{B}}$ for some special basis $\mathcal{B}$, we will conclude that $X$ is the high-ray space of some special order tree. To that aim, we will study suitable moves for Player $\mathrm{II}$ against a winning strategy for $\mathrm{II}$.
Before that, we will explicit a convenient method to partition a basic open set from a basis $\mathcal{B}$ generated by a special clopen subbase $\mathcal{C}=\{U_{\alpha}\}_{\alpha\in\Lambda}$. Considering the expression 5, this open set is written as $[U,F]=\displaystyle U\setminus\bigcup_{\alpha\in F}U_{\alpha}$, for some $U\in\mathcal{C}$ and some finite $F\subset\Lambda$. In particular, we can always assume that $U_{\alpha}\cap U_{\beta}=\emptyset$ for every pair $\alpha,\beta\in F$, because $\mathcal{C}$ is a nested family. By the same reason, $U_{\alpha}$ is a proper subset of $U$ for every $\alpha\in F$, unless $[U,F]=\emptyset$. When comparing basic open sets, the following observation is often implicitly applied:
Lemma 5.4.
Let $[U,F],[V,G]\in\mathcal{B}$ be basic open sets such that $[V,G]\subset[U,F]$. Then, $[V,G]=[W,K]$ for some subbasic clopen set $W\subseteq U$ and some finite $K\subset\Lambda$.
Proof.
Since $[V,G]\subset[U,F]$, we have $V\cap U\neq\emptyset$, unless $[V,G]=\emptyset$. If $V\subseteq U$, the result is immediate. If not, then $U\subseteq V$, because $\mathcal{C}$ is a nested family. In this case, we claim that $[V,G]=[U,F\cup G]$, finishing the proof. In fact, if $x\in[V,G]$, then $x\in[U,F]$ and $x\notin U_{\alpha}$ for any $\alpha\in G$. Hence, $x\in[U,F\cup G]$. Conversely, if $x\in[U,F\cup G]$, then, in particular, $x\notin U_{\alpha}$ for every $x\in V$ and any $\alpha\in G$, because $U\subseteq V$. In other words, $x\in[V,G]$. ∎
[…]
Lemma 5.5.
For a basic open set $[U,F]\in\mathcal{B}$, precisely one of the items below holds:
  • $i)$
    Either, for every $x\in[U,F]$, there is a clopen set $U_{x}\in\mathcal{C}$ such that $x\in U_{x}\subsetneq U$;
  • $ii)$
    Or there is a unique $x\in[U,F]$ such that $U$ is the smallest clopen set from $\mathcal{C}$ containing $x$, namely, there is no $V\in\mathcal{C}$ such that $x\in V\subsetneq U$.
Proof.
Assume that there are distinct $x,y\in[U,F]$ such that $U$ is the smallest clopen set from $\mathcal{C}$ containing them. Since $X$ is a Hausdorff topological space, there are $U_{x},U_{y}\in\mathcal{U}$ and finite sets $F_{x},F_{y}\subset\Lambda$ such that $x\in[U_{x},F_{x}]\subset[U,F]$ and $y\in[U_{y},F_{y}]\subset[U,F]$, but $[U_{x},F_{x}]\cap[U_{y},F_{y}]=\emptyset$. By the minimality of $U$, we must have $U\subseteq U_{x}$ and $U\subseteq U_{y}$, so that $x,y\in U\subseteq U_{x}\cap U_{y}$. We then can assume that $U_{x}\subseteq U_{y}$, because $\mathcal{C}$ is a nested family. Therefore, there is $i\in F_{y}$ such that $x\in U_{i}$, since $x\in U_{y}$ but $[U_{x},F_{x}]\cap[U_{y},F_{y}]=\emptyset$. Hence, $U\subseteq U_{i}$, again by the minimality of $U$. This, however, contradicts the fact that $y\in[U_{y},F_{y}]=\displaystyle U_{y}\setminus\bigcup_{j\in F_{y}}U_{j}$. ∎
Based on the cases displayed by Lemma 5.5, precisely one of the four items below describe a partition for $[U,F]$ into basic open sets from $\mathcal{B}$. For further reference, such an open cover will be denoted by $\mathcal{K}^{\prime}[U,F]$. In addition, although a basic open set $V\in\mathcal{K}^{\prime}[U,F]$ might admit more than one representation of the form $V=[U^{\prime},F^{\prime}]$, the following procedure will fix a clopen set $U^{\prime}\in\mathcal{C}$ and a finite $F^{\prime}\subset\Lambda$ such that $V=[U^{\prime},F^{\prime}]$. Moreover, this choice will be done so that $U^{\prime}\subseteq U$:
  1. 1.
    Assume first that item $i)$ from the above lemma is verified. Fix, for every $x\in[U,F]$, a clopen set $U_{x}\in\mathcal{C}$ such that $x\in U_{x}\subsetneq U$. If, for some $A\subset[U,F]$, the subfamily $\{U_{x}:x\in A\}$ is a chain regarding $\subseteq$, there is $\tilde{x}\in A$ such that $U_{\tilde{x}}$ is a maximum element in $\{U_{x}:x\in A\}$, because $\mathcal{C}$ is noetherian. Hence, by Zorn’s Lemma, the set
    $\tilde{A}=\{\tilde{x}\in[U,F]:U_{\tilde{x}}\text{ is }\subseteq-\text{maximal in }\{U_{x}\}_{x\in[U,F]}\}$
    is well-defined. Actually, this argument shows that, for every $x\in[U,F]$, there is $\tilde{x}\in\tilde{A}$ so that $x\in U_{\tilde{x}}$. Moreover, given $\tilde{x},\tilde{y}\in\tilde{A}$, then $U_{\tilde{x}}\cap U_{\tilde{y}}\neq\emptyset$ or $U_{\tilde{x}}=U_{\tilde{y}}$ by the $\subseteq$-⊆ -maximality of both $U_{\tilde{x}}$ and $U_{\tilde{y}}$, since $\mathcal{C}$ is nested. Hence, $\{U_{\tilde{x}}\}_{\tilde{x}\in\tilde{A}}$ is a disjoint family. In particular, for every $\alpha\in F$, there is a unique $\tilde{x}_{\alpha}\in\tilde{A}$ such that $U_{\alpha}\subset U_{\tilde{x}_{\alpha}}$. Then, let $U_{\alpha}^{\prime}$ be the $\subseteq$-⊆ -maximum set from $\{V\in\rho(U_{\alpha}):U_{\alpha}\subseteq V\subsetneq U_{\tilde{x}_{\alpha}}\}$. Within this notation,
    $\mathcal{K}^{\prime}[U,F]=\left\{U_{\tilde{x}}\setminus\bigcup_{\alpha\in F}U_{\alpha}^{\prime}:\tilde{x}\in\tilde{A}\right\}\cup\left\{U_{\alpha}^{\prime}\setminus\bigcup_{\begin{subarray}{c}\beta\in F\\ U_{\alpha}^{\prime}\not\subset U_{\beta}^{\prime}\end{subarray}}U_{\beta}^{\prime}:\alpha\in F\right\}$
    is a disjoint open cover for $[U,F]$ whose elements belong to the basis $\mathcal{B}$;
  2. 2.
    For this and the next two cases, assume that item $ii)$ of Lemma 5.5 holds. Moreover, for each $\alpha\in F$, denote by $U_{\alpha}^{\prime}$ the $\subseteq$-⊆ -maximum set from $\{W\in\rho(U_{\alpha}):U_{\alpha}\subsetneq W\subsetneq U\}$, if it exists. Otherwise, denote $U_{\alpha}^{\prime}=U_{\alpha}$. Define the set $F^{\prime}=\{\alpha\in F:U_{\alpha}^{\prime}\neq U_{\alpha}\}$. If $F^{\prime}\neq\emptyset$, then
    $\mathcal{K}^{\prime}[U,F]=\left\{U\setminus\bigcup_{\alpha\in F}U_{\alpha}^{\prime}\right\}\cup\left\{U_{\alpha}^{\prime}\setminus\bigcup_{\begin{subarray}{c}\beta\in F\\ U_{\alpha}^{\prime}\not\subset U_{\beta}^{\prime}\end{subarray}}U_{\beta}^{\prime}:i\in F^{\prime}\right\}$
    is a disjoint open cover for $[U,F]$ whose elements belong to the basis $\mathcal{B}$;
  3. 3.
    Following the notation from the above item, suppose now that $F^{\prime}=\emptyset$ and that $[U,F]=\{x\}$ is singleton. In this case, we set the trivial open cover $\mathcal{K}^{\prime}[U,F]=\{[U,F]\}=\{\{x\}\}$;
  4. 4.
    Finally, let $x$ be the unique point from $[U,F]$ such that $U$ is the smallest clopen set of $\mathcal{C}$ containing $x$. Following the notation from the previous items, suppose that $F^{\prime}=\emptyset$ and fix $y\in[U,F]\setminus\{x\}$. Then, there is $U_{y}\subsetneq U$ a subbasic set containing $y$. Since $y\notin U_{\alpha}$ for any $\alpha\in F$, it follows that $U_{y}\cap U_{\alpha}=\emptyset$, because $F^{\prime}=\emptyset$ and $\rho(U_{\alpha})$ is a cofinal sequence in $\mathring{\lceil U_{\alpha}\rceil}$. In this case,
    $\mathcal{K}^{\prime}[U,F]=\left\{U_{y},[U,F]\setminus U_{y}\right\}$
    is the claimed open cover for $[U,F]$.
Regarding these definitions, we can study convenient matches between Player $\mathrm{I}$ and a fixed winning strategy $\psi$ for Player $\mathrm{II}$. First, for a basic open set $[U,F]\in\mathcal{B}$, we remark that any other element from $[V,G]\in\psi([U,F])$ is contained in $[U,F]$. Hence, relying on Lemma 5.4, we can assume that $V\subseteq U$.
Now, suppose that Player $\mathrm{I}$ starts the game $\mathrm{End}_{\mathcal{B}}$ by declaring a basic open set $[U_{0},F_{0}]$. At the $n-$th round of the game, suppose that Player $\mathrm{II}$ declared a disjoint open cover $\mathcal{U}_{n}\subset\mathcal{B}$ for the basic open set $[U_{n},F_{n}]$ just played by Player $\mathrm{I}$. We write $\psi([U_{n},F_{n}])=\mathcal{U}_{n}$, assuming that the answer of Player $\mathrm{II}$ follows the given winning strategy. Then, we consider the case in which Player $\mathrm{I}$ starts the $(n+1)-$th round by choosing any basic open set $[U_{n+1},F_{n+1}]\in\displaystyle\bigcup_{[U,F]\in\mathcal{u}_{n}}\mathcal{K}^{\prime}[U,F]$. In particular, $[U_{n+1},F_{n+1}]\in\mathcal{K}^{\prime}[\tilde{U_{n}},\tilde{F_{n}}]$ for a unique basic open set $[\tilde{V_{n}},\tilde{F_{n}}]\in\mathcal{U}_{n}$. Therefore, we described a match
$\langle[U_{0},F_{0}],\mathcal{U}_{0},[U_{1},F_{1}],\mathcal{U}_{1},[U_{2},F_{2}],\mathcal{U}_{2},\dots\rangle.$
Since $\psi$ is a winning strategy for Player $\mathrm{II}$, we therefore can uniquely write
$\bigcap_{n\in\mathbb{N}}[U_{n},F_{n}]=\{x\}\cup A,$
where $A\subset X$ is open and $x\in X\setminus A$. Moreover, $[\tilde{U_{n+1}},\tilde{F_{n+1}}]\subset[U_{n+1},F_{n+1}]\subset[\tilde{U_{n}},\tilde{F_{n}}]$ for every $n\in\mathbb{N}$. Hence, although $A$ might not be a basic open set from $\mathcal{B}$, we will ensure below that $\{A\}\cup\mathcal{C}$ is a nested family. As a consequence, one can show that $A$ is a basic open set or a disjoint union of subbasic sets:
Proposition 5.6.
Given $U\in\mathcal{C}$, then $A\cap U=\emptyset$, $U\subseteq A$ or $A\subseteq U$.
Proof.
Suppose that $A\cap U\neq\emptyset$. In particular, $U\cap U_{n}\neq\emptyset$ for every $n\in\mathbb{N}$. If $U_{n}\subset U$ for some $n\in\mathbb{N}$, we are done, because $A\subset U_{n}$. Therefore, since $\mathcal{C}$ is a nested family, we can assume that $U\subsetneq U_{n}$ for each $n\in\mathbb{N}$. However, for a contradiction, suppose that $U\not\subset[U_{n},F_{n}]$ for some $n\in\mathbb{N}$. Hence, $U\cap U_{\alpha}\neq\emptyset$ for some $\alpha\in F_{n}$, so that $U_{\alpha}\subsetneq U$ (because $\mathcal{C}$ is nested and $U\cap A\neq\emptyset$). This verifies that the set
$\left\{U_{\beta}\in\mathcal{C}:\beta\in\displaystyle\bigcup_{m\in\mathbb{N}}U_{m},U_{\alpha}\subseteq U_{\beta}\subsetneq U\right\}$
is not empty and, since $\mathcal{C}$ is noetherian, there is $U_{\beta}$ a $\subseteq$-⊆ -maximal element. Hence, $\beta\in\displaystyle\bigcup_{m\geq n_{0}}U_{m}$ for some big enough $n_{0}\in\mathbb{N}$, since $[U_{m+1},F_{m+1}]\subset[U_{m},F_{m}]$ for every $m\in\mathbb{N}$. Besides that, if $m>n_{0}$, then $\mathcal{K}^{\prime}[\tilde{U_{m}},\tilde{F_{m}}]$ is not constructed as in items 3 and 4, because, once $U_{\beta}\subsetneq U\subsetneq\tilde{U_{m}}$, the set $\{V\in\rho(\tilde{U_{m}}):U_{\beta}\subsetneq V\subsetneq\tilde{U_{m}}\}$ is not empty.
Therefore, $\mathcal{K}^{\prime}[\tilde{U_{m}},\tilde{F_{m}}]$ is defined according to items 1 and 2. In particular, $U_{m+1}\in\rho(U_{\beta})$ for every $m>n_{0}$, because $U_{\beta}\subset U\subset\displaystyle\bigcap_{n\in\mathbb{N}}U_{n}$. This means that $\{U_{m}\}_{m>n_{0}+1}$ is an infinite subsequence of $\rho(U_{\beta})$, being also cofinal (regarding $\supseteq$) in $\mathring{\lceil U_{\beta}\rceil}$. Since $U_{\beta}\subsetneq U$, we must have $U_{m}\subsetneq U$ for some $m>n_{0}+1$, which is a contradiction.
Hence, we verified that $U\subset\displaystyle\bigcap_{n\in\mathbb{N}}U_{n}$. If $A\subset U$, we are done. If not, we will show that $x\notin U$, which concludes the inclusion $U\subset A$. To that aim, fix $y\in A\setminus U$ and suppose that $x\in U$. Since $U$ is a clopen set, $(A\setminus U)\setminus\{y\}$ is open, because so is $A\setminus U$. Then, we can write $\{x\}\cup A$ as $\{y\}\cup U\cup((A\setminus U)\setminus\{y\})$, contradicting the unique representation of $\bigcap_{n\in\mathbb{N}}[U_{n+1},F_{n+1}]$ as the disjoint union of a point and an open set. ∎
Corollary 5.7.
We can write $A=\displaystyle\bigcup_{U\in\mathcal{K}[A]}U$ for some disjoint family $\mathcal{K}[A]\subset\mathcal{C}$ or $A$ is a basic open set. If this latter case, we write $\mathcal{K}[A]:=A$
Proof.
First, suppose that there is $y\in A$ such that $A\subset U$ for every $U\in\mathcal{C}$ containing $y$. Since $A$ is open, there is $[U,F]\in\mathcal{B}$ such that $y\in[U,F]\subset A$. Hence, we must have $A\subset U$. Moreover, for every $\alpha\in F$ such that $A\cap U_{\alpha}\neq\emptyset$, we must have $U_{\alpha}\subseteq A$ by the above Proposition, because $y\notin U_{\alpha}$. Therefore, by considering the family $F^{\prime}=\{\beta\in F:A\cap U_{\beta}=\emptyset\}$, we conclude that $A=[U,F^{\prime}]$ is a basic open set.
Now, suppose that, for every $y\in A$, there is $U_{y}\in\mathcal{C}$ satisfying $y\in U_{y}\subset A$. Since $\mathcal{C}$ is nested and noetherian, we can apply Zorn’s Lemma to define the set
$\tilde{A}=\{\tilde{y}\in A:U_{\tilde{y}}\text{ is }\subseteq-\text{maximal in }\{U_{y}\}_{y\in A}\}$
and verify that $\mathcal{K}[A]:=\{U_{\tilde{y}}:\tilde{y}\in\tilde{A}\}$ is a disjoint cover for $A$. Actually, this is precisely the same argument presented by the definition of $\mathcal{K}^{\prime}[U,F]$ as in item 1. ∎
The description of the open set $A$ as in Corollary 5.7 suggests that the end game $\mathrm{End}_{\mathcal{B}}$ could be played again, with Player $\mathrm{I}$ starting the match by declaring $A$ (if it is a basic open set) or some clopen set from $\mathcal{K}[A]$. Hence, by iteratively playing $\mathrm{End}_{\mathcal{B}}$ against the winning strategy $\psi$ of Player $\mathrm{II}$, we can describe a tree $T_{\mathcal{C}}$ by induction as follows:
  • Without loss of generality, we can assume that $X\in\mathcal{C}$. Then, we set $X=[X,\emptyset]$ as the root of $T_{\mathcal{C}}$;
  • Every node of $T_{\mathcal{C}}$ is some basic open set $[U,F]$ from $\mathcal{B}$, possibly even a subbasic one when considering $F=\emptyset$. Moreover, $T_{\mathcal{C}}$ is ordered by the inverse inclusion $\supseteq$. In this case, we consider the (disjoint) open cover $\mathcal{W}=\psi([U,F])$. Then, the successors of $[U,F]$ in $T_{\mathcal{C}}$ are precisely the open sets from
    $\mathcal{K}[U,F]:=\displaystyle\bigcup_{[\tilde{U},\tilde{F}]\in\mathcal{W}}\mathcal{K}^{\prime}[\tilde{U},\tilde{F}],$
    described by the items 1-4;
  • For a fixed high-ray $R$ of $T_{\mathcal{C}}$, we consider $\{[U_{n},F_{n}]\}_{n\in\mathbb{N}}$ a cofinal sequence in $R$. We can see these open sets as answers for Player $\mathrm{I}$ in the match
    $\langle[U_{0},F_{0}],\psi([U_{0},F_{0}]),[U_{1},F_{1}],\psi([U_{1},F_{1}]),[U_{2},F_{2}],\psi([U_{2},F_{2}]),\dots\rangle$
    Since $\psi$ is a winning strategy for Player $\mathrm{II}$, we can write uniquely
    $\bigcap_{n=0}^{\infty}[U_{n},F_{n}]=\{x\}\cup A$
    for some point $x\in X$ and some open set $A\subset X$ not containing $x$. Then, we define $\mathcal{K}[A]$ as the set of tops of the high-ray $R$, where $\mathcal{K}[A]$ is given by Corollary 5.7. If $\{[V_{n},G_{n}]\}_{n\in\mathbb{N}}$ is another cofinal sequence in $R$, then $\bigcap_{n\in\mathbb{N}}[U_{n},F_{n}]=\displaystyle\bigcap_{n\in\mathbb{N}}[V_{n},G_{n}]$, because $[U_{n+1},F_{n+1}]\subset[U_{n},F_{n}]$ and $[V_{n+1},G_{n+1}]\subset[V_{n},G_{n}]$ for every $n\in\mathbb{N}$. In other words, the description of $A$ does not depend on the choice of $\{[U_{n},F_{n}]\}_{n\in\mathbb{N}}$. Hence, this procedure defines the limit nodes of $T_{\mathcal{C}}$.
Relying on the fact that $\mathcal{C}$ is $\sigma-$disjoint, we will now argue how $T_{\mathcal{C}}$ is a special tree. To this aim, we observe that, for every $U\in\mathcal{C}$, it is enough to construct a partition $\{\mathcal{A}_{U}^{n}\}_{n\in\mathbb{N}}$ of
$T_{U}=\{[U,F]:[U,F]\in T_{\mathcal{C}}\text{ for some finite }F\subset\Lambda\}$
into countably many disjoint families. This because, if $\{\mathcal{C}_{k}\}_{k\in\mathbb{N}}$ is a partition of $\mathcal{C}$ into antichains, then $\{T_{k}^{n}:n,k\in\mathbb{N}\}$ is a partition of $T$ into antichains, where $T_{k}^{n}=\displaystyle\bigcup_{U\in\mathcal{C}_{k}}\mathcal{A}_{U}^{n}$ for every $n,k\in\mathbb{N}$. Indeed, the description of $\{\mathcal{A}_{U}^{n}\}_{n\in\mathbb{N}}$ is the core of the following lemma:
Lemma 5.8.
For each $U\in\mathcal{C}$, there is $\{\mathcal{A}_{U}^{n}\}_{n\in\mathbb{N}}$ a cover of $T_{U}$ by antichains. In particular, $T_{\mathcal{C}}$ is a special tree.
Proof.
Consider $\mathcal{A}_{U}^{0}$ the set of $\subseteq$-⊆ -maximal elements from $T_{U}$, or, equivalently, the set of minimal elements of $T_{U}$ is the tree order of $T_{\mathcal{C}}$. Clearly, $\mathcal{A}_{U}^{0}$ is an antichain. For each $V\in\mathcal{A}_{U}^{0}$, it is enough to conclude that the set $\downarrow V=\{V^{\prime}\in T_{U}:V^{\prime}\subsetneq V\}$ is countable, written as $\{V_{i}\}_{i<\kappa_{V}}$ for some cardinal $\kappa_{V}\leq\omega$. Then, for $n\geq 1$, the claimed antichain $\mathcal{A}_{U}^{n}$ can be given by $\mathcal{A}_{U}^{n}=\{V_{n}:V\in\mathcal{A}_{U}^{0}\text{ and }\kappa_{V}\geq n\}$.
In fact, a basic open set $V\in\mathcal{A}_{U}^{0}$ can be written as $[U,F]$ for some finite $F\subset\Lambda$. First, suppose that, for every $x\in V$, there is a clopen set $U_{x}\in\mathcal{C}$ such that $x\in U_{x}\subsetneq U$. For a basic open set $[U_{0},F_{0}]\in\psi([U,F])$, we have either $U_{0}\subsetneq U$ or $U_{0}=U$. In the former case, every element from $\mathcal{K}^{\prime}[U_{0},F_{0}]$ can be written as $[U_{1},F_{1}]$ for some clopen set $U_{1}\subseteq U_{0}\subsetneq U$. In the latter, $\mathcal{K}^{\prime}[U_{0},F_{0}]$ is defined according to item 1, by the assumption over $U$. Then, every element from $\mathcal{K}^{\prime}[U_{0},F_{0}]$ has the form $[U_{1},F_{1}]$ for some clopen set $U_{1}\subsetneq U_{0}=U$. In both cases, $U_{1}\subsetneq U$ for each $[U_{1},F_{1}]\in\mathcal{K}[U,F]$. As a consequence, for every $[U_{2},F_{2}]\in T_{\mathcal{C}}$ such that $[U_{2},F_{2}]\subsetneq[U,F]$, we must have $U_{2}\subsetneq U$. Hence, $\downarrow V=\emptyset$ in this case.
Now, according to the dichotomy of Lemma 5.5, we assume the existence of a unique $x\in[U,F]$ such that $U$ is the smallest clopen set from $\mathcal{C}$ containing $x$. Then, $\mathcal{K}^{\prime}[\tilde{U},\tilde{F}]$ is constructed following items 2-4, if $[\tilde{U},\tilde{F}]$ denotes the open set from $\psi([U,F])$ containing $x$. In any of the three criteria, there is a unique $V^{\prime}\in\mathcal{K}[U,F]$ of the form $V^{\prime}=[U,F_{1}]$ for some finite $F_{1}\subset\Lambda$, i.e., $T_{U}\cap\mathcal{K}[U,F]=\{V^{\prime}\}$. Indeed, $V^{\prime}$ is the open set from $\mathcal{K}[U,F]$ containing $x$. By induction, suppose that we have defined $F_{1},F_{2},\dots,F_{n}\subset\Lambda$ such that $T_{U}\cap\mathcal{K}[U,F_{i}]=\{[U,F_{i+1}]\}$ for every $1\leq i<n$. Moreover, assume that $x\in[U,F_{i}]$ for every $1\leq i\leq n$. As before, item $ii)$ of Lemma 5.5 holds, so that the open set from $\mathcal{K}[U,F_{n}]$ containing $x$ is the unique that can be written as $[U,F_{n+1}]$ for some finite $F_{n+1}\subset\Lambda$. More precisely, $T_{U}\cap\mathcal{K}[U,F_{n}]=\{F_{n+1}\}$, because, if $[\tilde{U},\tilde{F}]$ denotes the open set from $\psi([U,F])$ containing $x$, the partition $\mathcal{K}^{\prime}[\tilde{U},\tilde{F}]$ was again constructed according to items 2-4.
At the end of this recursive process, $\{[U,F_{n}]\}_{n\in\mathbb{N}}$ defines a cofinal sequence in a high-ray $R$. Alternatively, we can see $\{[U,F_{n}]\}_{n\in\mathbb{N}}$ as the moves of Player $I$ in the match
$\langle[U,F_{0}],\psi([U,F_{0}]),[U,F_{1}],\psi([U,F_{1}]),[U,F_{2}],\psi([U,F_{2}]),\dots\rangle$
For instance, suppose that there is an open neighborhood of $x$ in $\bigcap_{n=0}^{\infty}[U,F_{n}]$ containing at least two elements. Since $U$ is the smallest clopen set from $\mathcal{C}$ containing $x$, we can represent this open neighborhood as $[U,F_{\infty}]$ for some finite $F_{\infty}\subset\Lambda$. Assuming that $F_{\infty}$ is $\subseteq$-⊆ -minimal with this property, for every $\alpha\in F_{\infty}$ there are $n_{\alpha}\in\mathbb{N}$ and $\eta_{\alpha}\in F_{n_{\alpha}}$ such that $U_{\eta_{\alpha}}\subseteq U_{\alpha}$. If $U_{\eta_{\alpha}}$ is a proper subset of $U_{\alpha}$, then $\{W\in\rho(U_{\eta_{\alpha}}):U_{\eta_{\alpha}}\subsetneq W\subsetneq U\}$ is not empty, so that $\mathcal{K}^{\prime}[\tilde{U},\tilde{F}]$ is constructed as in item 2. Again, $[\tilde{U},\tilde{F}]\in\psi([U,F_{n_{\alpha}}])$ denotes the open set containing $[U,F_{n_{\alpha}+1}]$. Hence, there is $\eta_{\alpha}^{\prime}\in F_{n_{\alpha}+1}$ such that $U_{\eta_{\alpha}}\subsetneq U_{\eta_{\alpha}^{\prime}}\subseteq U_{\alpha}$. Since there are no infinite $\subseteq$-⊆ -increasing chains in $\mathcal{C}$, we can actually choose, for each $\alpha\in F_{\infty}$, a index $n_{\alpha}\in\mathbb{N}$ so that $\alpha\in F_{n_{\alpha}}$.
Therefore, if $n=\max\{n_{\alpha}:\alpha\in F_{\infty}\}$, then $\mathcal{K}^{\prime}[\tilde{U},\tilde{F}]$ is described by item 4, where $[\tilde{U},\tilde{F}]\in\psi([U,F_{n}])$ is the open set containing $[U,F_{n+1}]$. Then, $F_{n+1}=F_{n}\cup\{\beta\}$ for some index $\beta\in\Lambda\setminus F_{n}$ such that $U_{\beta}\cap U_{\beta^{\prime}}=\emptyset$ for every $\beta^{\prime}\in F_{n}$. On the other hand, $F_{n}=F_{\infty}$ by the choice of $n$ and the $\subseteq$-⊆ -minimality of $F_{\infty}$. Hence, we contradict the fact that $[U,F_{\infty}]\subset[U,F_{n+1}]$.
In other words, we proved that, when writing $\bigcap_{n=0}^{\infty}[U,F_{n}]$ as a disjoint union of a point and an open set $A$, this point must be $x$. By Proposition 5.7, $A$ has either the form $[W,H]\in\mathcal{B}$ or it is a disjoint union of a family $\mathcal{K}[A]\subset\mathcal{C}$. In the former case, we note that $W\subsetneq U$: otherwise, $W=U$ and $x\in U_{\alpha}$ for some $\alpha\in H$, contradicting the minimality of $U$ as a clopen set from $\mathcal{C}$ containing $x$. In the latter, $U\notin\mathcal{K}[A]$ because $x\notin A$. In both cases, if $[U^{\prime},F^{\prime}]\in T_{\mathcal{C}}$ is contained in some top of $R$, then $U^{\prime}\subsetneq U$. This proves that $\downarrow V=\{[U,F_{n}]:n\geq 1\}$. ∎
In particular, the last paragraph of the above proof verifies the following useful property:
Corollary 5.9.
Consider an infinite chain in $T_{\mathcal{C}}$ of the form $\{[U,F_{n}]\}_{n\in\mathbb{N}}$, for some $U\in\mathcal{C}$. Then, let
$\bigcap_{n=0}^{\infty}[U,F_{n}]=\{x\}\cup A$
be the unique description of $\bigcap_{n=0}^{\infty}[U,F_{n}]$ as the disjoint union of a point and an open set. Then, as in item $ii)$ of Lemma 5.5, $U$ is the smallest clopen set of $\mathcal{C}$ containing $x$.
To summarize, $T_{\mathcal{C}}$ displays suitable basic open sets from $\mathcal{B}$ via a tree ordered by $\supseteq$. In particular, any chain in $T_{\mathcal{C}}$ of the form $[U_{0},F_{0}]\subseteq[U_{1},F_{1}]\subseteq[U_{2},F_{2}]\subseteq\dots$ must stabilize, i.e., there must exist $[U_{n},F_{n}]$ a $\subseteq$-⊆ -maximal element for some $n\in\mathbb{N}$. Moreover, if $[U,F]$ and $[V,G]$ are incomparable in the tree order of $T_{\mathcal{C}}$, so that $[U,F]\not\subseteq[V,G]$ and $[V,G]\not\subseteq[U,F]$, then $[U,F]\cap[V,G]=\emptyset$. This because $\bigcap\{[W,H]\in T_{\mathcal{C}}:[W,H]\subseteq[U,F]\text{ and }[W,H]\subseteq[V,G]\}$ is either a basic open set $I=[W^{\prime},H^{\prime}]$ or a disjoint union of a point and an open set $I$. In both cases, $[U,F]$ and $[V,G]$ are subsets of distinct elements from a disjoint open cover for $I$. In particular, any antichain in the tree $T_{\mathcal{C}}$ is an antichain of $X$ as a topological space. In other words, the set $\psi(\mathcal{C}):=\{[U,F]\in\mathcal{B}:[U,F]\in T_{\mathcal{C}}\}$ of nodes of $T_{\mathcal{C}}$ is a noetherian, nested and $\sigma-$disjoint family. Moreover, its elements are also clopen sets, as observed by Lemma 5.1.
In particular, for every $x\in X$, the set $\mathcal{V}_{x}^{\prime}=\{[U,F]\in\psi(\mathcal{C}):x\in[U,F]\}$ is a chain in $\psi(\mathcal{C})$, because, once $\psi(\mathcal{C})$ is nested, its elements are pairwise comparable. Besides that, for every $[U,F]\in\mathcal{V}_{x}^{\prime}$, there is a (unique) basic open set from $\mathcal{K}[U,F]$ containing $x$, since $\mathcal{K}[U,F]$ is a disjoint open cover for $[U,F]$. Therefore, $\mathcal{V}_{x}^{\prime}$ describes a high-ray in $T_{\mathcal{C}}$, so that $\bigcap\mathcal{V}_{x}^{\prime}$ is written uniquely as a union of a point and an open set $A$ not containing it. We observe that this point is precisely $x$: otherwise, $x$ belongs to some $[U,F]\in\mathcal{K}[A]$, contradicting the definition of $\mathcal{V}_{x}^{\prime}$. Due to this fact, we write $A_{x}=A$ and state the following remark:
Proposition 5.10.
For every $x\in X$, fix $\{[U_{n},F_{n}]\}_{n\in\mathbb{N}}$ a countable cofinal sequence in $\mathcal{V}_{x}^{\prime}$ regarding the tree order of $T_{\mathcal{C}}$, whose existence follows from the fact that $T_{\mathcal{C}}$ is special. Then, the set
$\mathcal{V}_{x}=\left\{[U_{n},F_{n}]\setminus\bigcup_{U\in\mathcal{F}}U:n\in\mathbb{N},\mathcal{F}\subset\mathcal{K}[A_{x}]\text{ finite}\right\}$
is a local basis for $x$. In particular, by Lemma 5.1, $\psi(\mathcal{C})$ is a special clopen subbase for $X$.
Proof.
Since $\{[U_{n},F_{n}]\}_{n\in\mathbb{N}}$ is cofinal in $\mathcal{V}_{x}^{\prime}$, we have $\bigcap_{n=0}^{\infty}[U_{n},F_{n}]=\bigcap_{A\in\mathcal{V}_{x}^{\prime}}A=\{x\}\cup A_{x}$. Then, let $[W,H]$ be a basic open set from $\mathcal{B}$ containing $x$. For every $\theta\in H$, we can assume that $U_{\theta}\cap[U_{n},F_{n}]\neq\emptyset$ for every $n\in\mathbb{N}$: in this case, an open set $V\in\mathcal{V}_{x}$ such that $x\in V\subset[W,H\setminus\{\theta\}]$ also satisfies $x\in V\subset[W,H]$. In particular, $U_{n}\cap U_{\theta}\neq\emptyset$ for every $n\in\mathbb{N}$, so that $U_{\theta}\subsetneq U_{n}$ for every $n\in\mathbb{N}$ because $\mathcal{C}$ is nested and $x\in U_{n}\setminus U_{\theta}$. Analogously, for every $\theta\in H$ and every $\alpha\in\displaystyle\bigcup_{n\in\mathbb{N}}F_{n}$, we can assume that $U_{\theta}\cap U_{\alpha}=\emptyset$ or $U_{\alpha}\subsetneq U_{\theta}$. Otherwise, since $\mathcal{C}$ is nested, we would have $U_{\theta}\subseteq U_{\alpha}$ for some $\alpha\in F_{n}$ and some $n\in\mathbb{N}$. Actually, there would exist, for every $k\geq n$, an index $\alpha^{\prime}\in F_{k}$ such that $U_{\theta}\subseteq U_{\alpha^{\prime}}\subseteq U_{\alpha}$. In this case, if $V\in\mathcal{V}_{x}$ satisfies $x\in V\subset[W,H]$ and has the form $[U_{m},F_{m}]\setminus\bigcup_{U\in\mathcal{F}}U$ for some $m\in\mathbb{N}$, then $V^{\prime}=\displaystyle[U_{m+n+1},F_{m+n+1}]\setminus\bigcup_{U\in\mathcal{F}}U$ is an open set from $\mathcal{V}_{x}$ such that $V^{\prime}\subset[W,H\setminus\{\theta\}]$.
Within the above considerations, suppose first that $U_{\alpha}\subsetneq U_{\theta}$ for some $\alpha\in\displaystyle\bigcup_{n\in\mathbb{N}}F_{n}$ and some $\theta\in H$. We can choose $\alpha$ so that $U_{\alpha}$ is $\subseteq$-⊆ -maximal satisfying $U_{\alpha}\subsetneq U_{\theta}$, because $\mathcal{C}$ is noetherian. If $n\in\mathbb{N}$ is an index such that $\alpha\in F_{n}$, then $U_{\alpha}\subsetneq U_{\theta}\subsetneq U_{n}$. However, if $[\tilde{U}_{n},\tilde{F}_{n}]\in\psi([U_{n},F_{n}])$ is the open set containing $[U_{n+1},F_{n+1}]$, then $\mathcal{K}^{\prime}[\tilde{U}_{n},\tilde{F}_{n}]$ is constructed following items 1 or 2. In both cases, there is $\alpha^{\prime}\in F_{n+1}$ such that $U_{\alpha}\subsetneq U_{\alpha^{\prime}}$. By the choice of $\alpha$, we must have $U_{\theta}\subseteq U_{\alpha^{\prime}}$, a situation that was discarded by the previous paragraph. Therefore, from now on, we assume that $U_{\theta}\cap U_{\alpha}=\emptyset$ for every $\theta\in H$ and every $\alpha\in\displaystyle\bigcup_{n\in\mathbb{N}}F_{n}$.
On the other hand, suppose that $W\subsetneq U_{n}$ for every $n\in\mathbb{N}$. Then, there must exist $n_{0}\in\mathbb{N}$ such that $[W,H]\cap U_{\alpha}\neq\emptyset$ for some $\alpha\in F_{n_{0}}$: otherwise, $x\in[W,F]\subset\bigcap_{A\in\mathcal{V}_{x}^{\prime}}A$, contradicting the uniqueness of the representation $\bigcap_{A\in\mathcal{V}_{x}^{\prime}}A=\{x\}\cup A_{x}$. Then, $U_{\alpha}\subsetneq W$, because $\mathcal{C}$ is a nested family and $x\in W\setminus U_{\alpha}$. Moreover, since $\mathcal{C}$ is noetherian, we assume that $U_{\alpha}$ is a $\subseteq$-⊆ -maximal element from $\bigcup_{n=1}^{\infty}F_{n}$ such that $U_{\alpha}\subsetneq W$. Therefore, one of the following cases is verified, but both lead to contradictions:
  • Suppose for instance that $U_{n+1}\subsetneq U_{n}$ for each $n\in\mathbb{N}$. By the $\subseteq$-⊆ -maximality of $U_{\alpha}$, we have $\alpha\in F_{n}$ for every $n\geq n_{0}$. Regarding that $U_{\alpha}\subsetneq W\subsetneq U_{n}$, the family $\{U_{n}\}_{n\in\mathbb{N}}$ is cofinal in $\mathring{\lceil U_{\alpha}\rceil}$. After all, for any $n\geq n_{0}$, the partition $\mathcal{K}^{\prime}[\tilde{U}_{n},\tilde{F}_{n}]$ was defined according to items 1 or 2, where $[\tilde{U_{n}},\tilde{F_{n}}]\in\psi([U_{n},F_{n}])$ is the basic open set containing $[U_{n+1},F_{n+1}]$. Then, we must have $U_{n}\subseteq W\subsetneq U_{n}$ for some big enough $n\geq n_{0}$, which is a contradiction;
  • If the above item does not hold, we can assume that $n_{0}\in\mathbb{N}$ is big enough in order to the equality $U_{n}=U_{n_{0}}$ be verified for every $n\geq n_{0}$. In this case, we are under the hypothesis of Corollary 5.9, so that $U_{n_{0}}$ is the smallest clopen set from $\mathcal{C}$ containing $x$. However, this contradicts the assumption that $x\in W\subsetneq U_{n_{0}}$.
To summarize, there must exist $n\in\mathbb{N}$ such that $U_{n}\subseteq W$. On the other hand, given $\theta\in H$, we remarked in the first two paragraphs that $U_{\theta}\subsetneq U_{m}$ for every $m\in\mathbb{N}$ and $U_{\theta}\cap U_{\alpha}=\emptyset$ for every $\alpha\in\displaystyle\bigcup_{m\in\mathbb{N}}F_{m}$. In other words, we proved that $U_{\theta}\subset\displaystyle\bigcap_{m\in\mathbb{N}}[U_{m},F_{m}]$. Hence, $U_{\theta}\subset A_{x}$ for every $\theta\in H$, because $x\notin U_{\theta}$. Since $\mathcal{K}[A_{x}]$ is singleton or a disjoint family of elements from $\mathcal{C}$, we observe that =$\mathcal{F}=\{B\in\mathcal{K}[A_{x}]:U_{\theta}\subset B\neq\emptyset\text{ for some }\theta\in H\}$ is finite. Hence, $[U_{n},F_{n}]\setminus\displaystyle\bigcup_{U\in\mathcal{F}}U\in\mathcal{V}_{x}$ and $x\in[U_{n},F_{n}]\setminus\bigcup_{U\in\mathcal{F}}U\subset[W,H]$, finishing the proof.
Therefore, combining Proposition 5.10 and Lemma 5.1, we see that $\psi(\mathcal{C})$ is a special clopen subbase for $X$. We will finish the proof that $X$ is the end space of a special tree, or, equivalently, the end space of some graph, via the recent characterization given by Pitz in [15]. His main result claims that these spaces are precisely the ones that admit a $\sigma-$disjoint, nested, noetherian and hereditarily complete clopen subbase.
In order to introduce this additional property, we recall that a family of sets $\mathcal{S}$ is centered if $\bigcap_{F\in\mathcal{F}}F\neq\emptyset$ for every finite $\mathcal{F}\subset\mathcal{S}$. Then, we say that $\mathcal{S}$ is complete if $\bigcap_{F\in\mathcal{F}}F\neq\emptyset$ for every centered family $\mathcal{F}\subset\mathcal{S}$. Finally, when $\mathcal{S}$ is a subfamily of the power set $\wp(X)$ of the topological space $X$, we call $\mathcal{S}$ hereditarily complete if $\{S\cap Y:S\in\mathcal{S}\}$ is complete for every closed subspace $Y\subset X$. Relying on Proposition 5.10, its not hard to conclude that $\psi(\mathcal{C})$ has this property:
Theorem 5.11.
The family $\psi(\mathcal{C})$ of the nodes of the tree $T_{\mathcal{C}}$ is hereditarily complete.
Proof.
Let $Y\subset X$ be a closed subspace and fix $\mathcal{S}_{Y}\subset\{Y\cap[U,F]:[U,F]\in\psi(\mathcal{C})\}$ a centered family. Then, $(Y\cap[W,H])\cap(Y\cap[V,G])\neq\emptyset$ for any two open sets $[W,H],[V,G]\in\mathcal{S}=\{[U,F]\in\psi(\mathcal{C}):[U,F]\cap Y\in\mathcal{S}_{Y}\}$. In particular, since $\psi(\mathcal{C})$ is nested, $\mathcal{S}$ is a chain in $T_{\mathcal{C}}$. If its cofinality is finite, then $\mathcal{S}$ has a $\subseteq$-⊆ -minimum element $[U,F]$, so that $\bigcap_{S\in\mathcal{S}}S=[U,F]\cap Y\neq\emptyset$.
Assume now that $\mathcal{S}$ has infinite cofinality in $T$, so that $\bigcap_{S\in\mathcal{S}}S$ is written uniquely as a disjoint union $\{x\}\cup A_{x}$ for some $x\in X$. For a contradiction, suppose that ($(\{x\}\cup A_{x})\cap Y=\emptyset$. Therefore, since $X\setminus Y$ is open, by Proposition 5.10 there are $[U,F]\in\mathcal{S}$ and a finite family $\mathcal{F}\subset\mathcal{K}[A_{x}]$ such that $x\in[U,F]\setminus\bigcup_{F\in\mathcal{F}}F\subset X\setminus Y$. Hence, $[U,F]\subset X\setminus Y$, because $\bigcup_{F\in\mathcal{F}}F\subset A_{x}$. This, however, contradicts the definition of $S$. ∎
In other words, Theorem 5.11 finishes a description of topological properties of a suitable clopen subbase $\psi(\mathcal{C})$ for $X$. More precisely, our study carried out in this section concludes that, if $X$ is a Hausdorff space in which $\mathrm{II}\uparrow\mathrm{End}_{\mathcal{B}}$ for some basis $\mathcal{B}$ generated by a given special clopen subbase $\mathcal{C}$, then there is $\psi(\mathcal{C})$ another special clopen subbase for $X$ that is hereditarily complete. According to Theorem $1.2$ of Pitz in [15], this proves that $X$ is the end space of a special tree. As formalized by Proposition 5.3, Player $\mathrm{II}$ also has a winning strategy for the end game in these latter spaces, which, by Theorem $3$ of Kurkofka and Pitz in [13], are precisely the topological spaces that arise as end spaces of graphs. Therefore, Theorem 1.2 is established.
6 Acknowledgments
The first and the third named authors thank the support of Fundação de Amparo à Pesquisa do Estado de São Paulo (FAPESP), being sponsored through grant numbers 2023/00595-6 and 2021/13373-6 respectively. In its turn, the second named author acknowledges the support of Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq) through grant number 165761/2021-0.